A Bibliography of Publications in SIAM Journal on Financial Mathematics

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

17 October 2023 Version 1.26

Title word cross-reference

1 [CCL^{+}11]. 3/2 [Liu15]. α [CMP21]. CARMA(p, q) [MPR21]. K [Shi17]. Λ [BP22]. H^{3} [FZ16]. t [KP18]. X [CSS20].

-Hypergeometric [CMP21]. -Quantiles [BP22]. -Scheme [Shi17]. -Value [CSS20].

500 [Guy22].

Ability [Cot21]. Absence [AS10]. Accidents [CF15]. Accuracy [FHS22]. Accurate [War13]. Across [LLP11]. Adapted [BW23]. Adaptive [CL21a, NV22]. Adjoint [PPR13]. Adjusted [BFZ21]. Adjustment [AC17].

Adjustments [BPO18, CCFK20, CSS20, dRP22]. Advantageous [For20]. Affect [DJ23]. Affine [CLP18, CW12, EGG10, Fon23, GG14, GPSS15, JMP21, JO23]. Agency [MR18]. Agent [BH16, FT22, HZ16, MR18, Zha23]. Agents [Lon20]. Algorithm [FJO21, HLLR16, LLP11, PPS18]. Algorithmic [CDJ17]. Algorithms [BBH^{+}21, BDL11, SZH13]. Allocation [ACDP18, For20, PBH10]. Almgren [ANW23]. Almgren-Chriss [ANW23]. Alpha [LLX19]. Alpha-Maxmin [LLX19]. Alternative [Fon23, vSDF19]. Always [Zha23]. Ambiguity [PW22]. American [ATL12, BZ14, BHZ15, BL16, CDP21, CPZ22, DK18, GL22, GHK21, How12, HRW13, JSDN11, JV11, LT19, Lel18, LHJ10, Liu15, PPS18, RW12, Zho21]. Ameripean [JSDN11]. Analysis [ALMY23, BH16, BBH^{+}21, Bic12, BGH21, BK18, CY21, FHS22, GV15, HK18, LLM^{+}23, LS21, Mon13a, NW21, RS12, SZH13, SSZ16, ZL22]. Analytic [FL13a, KP16]. Analytical [FT20, GP15]. Analytically [CFR13]. Anchor [GK22]. Anomalous [JT20]. Anticipated [CSS20]. Application [CM12, EM21, HO11, SZH13, Shi17, War13]. Applications [BZ14, CCL^{+}11, FL13a, GKR14, HLLR16, KP16, Li22, Str14]. Approach [BCN23, BAGM20, BHY19, BK15, BGO21, BDL11, CG20, CCY12, COW19, COW22, CL21a, CK18, EJJ15, FMM11a, FHS22, HS12, LLP11, LLL21, LMS22, MW13, PW22, Pun21, SZ21, VFL23]. Approaches [FKV12]. Approximating [BEV14]. Approximation [AE19, BDN15, BST10, BTW11, CFP10, FR14, GS12, GKS20, Guy22, HRW13, KM23, KN18, Lev11, MPR21, NZ13, PVW17, Yam22, ZL22]. Approximations [BDG18, CCL^{+}11, GP15, How12, HO11, JT11b, LS16, RBG21, VG22]. Arbitrage [BDG22, Bur16, CT15, GJMN16, JT11a, Lep16, MM22, NJZB23, Str14]. Arbitrage-Free [GJMN16, NJZB23]. ArgMax [KHOL10]. Arising [DB20, FJO21, IR21]. Arithmetic [Lev18]. Asian [FMM11b, How12, Kir16, Kol15, Lev18, PZ16, ZO13]. Ask [ZRA14]. Asset [BMMB18, BMMB19, DL11a, EV16, For20, JKP11, RBG21, RS12, BDG22]. Assets [EGLO21, FT20, JV11, Lep16, LPY21, LLCMA16]. Asymmetric [GL22]. Asymptotic [ALMY23, Bic12, CY21, GJRS18, Gul10, GV15, How12, JT11b, KN18, LX15, Mon13b, SSZ16, TY12]. Asymptotics [AFLZ17, CC16, CCSW19, CCL^{+}11, EFGR19, FP15, FZ16, FZ17, FHS22, JR15, FFF10]. At-the-Money [EFGR19]. Autoencoders [NJZB23].

Average [BTW11, CJ16]. Averaging [How12]. Averse [DV15, HZ16]. Aversion [DM22, FP15, MW20]. Aware [JPWT22]. Axiom [FLW23]. Axiomatization [BP22, FLW23].

Back [KHOL10]. Backward

[BQY22, BCC15, BBC16, CSS20, CM12, Lev18, MT17]. Balanced [QC19]. Bang [AF15]. Bang-Bang [AF15]. Bank [LLM^{+}23]. Banking [BMMB19]. Bankrupt [LLS14]. Bankruptcy [BMNP17, CCSW19, LPY21]. Barrier [ACN13, CO11, DL11b, FO11, Lev11, ZX19]. Based [AJ15, BAGM20, BHU18, CGSF23, FO11, HK18, JS11, ZO13]. Basis [AKK13, BHSW15, CLP11, LR16]. Basket [FZ16, HSZ17]. Bayesian [CK18, MP23]. Be [For20, VFL23, WBT15]. Beating [VFL23]. Behavior [AJ15, GJRS18]. Behavioral [CD12]. Beliefs [MWZZ22]. Benchmark [VFL23, Wan22]. Benchmarking [AS18]. Benefit [HZK17]. Bequest [BPY16]. Bergomi [Guy22]. Bermudan [BEST23, BDN15, BBC16, FO11, FJJ20, FL13b, Rog10, SZH13]. Best [BP13]. Between [ARS21, JSDN11, NM13]. Beyond [CCC^{+}21]. Bias [ATL12, GPS22]. Bid [ZRA14]. Bidding [BF14b]. Bilateral [LSZ20]. Binding [LMK13]. binomial [LHJ10]. Bivariate [FZ16]. Black [BHSW15, GS17, Käl20, Teh16]. Bond [CL21b]. Book [AJ15, AKU21, AS10, ASS12, CM21, HK17b, HR17, JL18, KM23, PSS11]. Books [HN14, HX19]. Border [CFVS22]. Borrowing [BF19]. boundaries [LHJ10]. Boundary [CL21b, JV11]. Bounded [CC16]. Bounds [Ara11, BND20, FN18, NS21, Pap18, Teh16]. Bridge [JSDN11]. Brokerage [ANW23]. Brownian [AAM22, FÖ21, GNR19]. BSDE [CG20, LZ17, MW13]. BSDEs [GP15, KP16, LR16, YZZ23]. BSPDEs [Li22]. Bubble [JKP11, SS19]. Bubbles [BN15, BMMB18, BMMB19, EJM21]. Buy [CJR14].

Càdlàg [Ara14, LR16]. Calculus [ATL12, Mon13b]. Calibrated [CC12]. Calibration [BLP16, CDK10, CMR19, CGSF23, DT13, GPSS15, GLOW22, Guy22]. Call [FLGL18, Gul10, MS10]. Calls [Fei22]. Can [WBT15]. Capital [BHY19, CSS20, LPY21, MW20, NZ19, Zhi23]. Caplet [Fon23]. Capped [BL13]. Caps [DK18]. CARA [MWZZ22]. Carlo [AN15, BHU18, GR12, MPR14, ZX19]. Carr [Lev11]. Case [ABY23, BZZ19, BW11, LSWY18]. Cash [PVW17]. Cash-Constrained [PVW17]. Category [Arm18]. CDO [CDK10, GS12, HO11]. Central [AFM20, AC17]. Centrality [BBC^{+}20]. Certain [JL13, KP10a]. Certainty [VRT22]. Chain [ZL22]. Chains [DR18]. Chances [EE22]. Change [HKMR20]. Changed [CLL21]. Changes [LLCMA16]. Chaos [FÖ21, Lel18]. Characteristic [FL13a]. Characteristics [JL20b]. Characterization [BST10, CD13, JL20a, Qi23]. Chebyshev [GKS20]. Chief [CS10]. Choice [BGZ22, MZ10]. Chriss [ANW23]. CIR [BZ20]. Claim [HX16]. Claims [AM11, PP10]. Class [BMZ11, CNP12, GR12]. Classes [NW21]. Clawback [SW22]. Clearing [AC17, FPR^{+}18, Fei22, KV19]. Clients [ANW23]. Climate [BGH21]. Climate-Economic [BGH21]. Cliquet [BL13]. Closed [CJ16, CCL^{+}11]. Closed-Form [CJ16, CCL^{+}11]. Clustering [SS15]. Coefficient [LS16]. Coefficients [CJM16, HSX23, ZL22]. Coherent [DR18]. Cointegration [CCW19]. Collateralization [BR15]. Collocation [DB20]. Commoditie [GTW19].

Commodity [CJQ16, CSS15, GTW19, HF10, JS11]. Common [GK10, MR18]. Communication [AG20, ACEL23, BD21, BFN22, BCDD21, BWZ22, BP22, BC22, BGH21, BFZ21, CCC^{+}21, DP20, DM22, DZ23, EJM21, EE22, FS21, Fei22, Fon23, FJO21, GHK21, MR18, Sap20, SZ22b, TD22, WXXY22, Yam22, Zha23]. Compactness [VF16]. Comparisons [Pap18]. Competition [AGZ22, HMKSY20, Lud11]. Complete [BDH10, HKZ17, WX21]. Complexity [BHU18]. Computation [BST10, BPO18, HSX15]. Computationally [MGH18]. Concerns [BLD17]. Conditional [ACN13, BH16, DF21, FKV12, For20, FM11]. Conditioned [DF18]. Conditioning [MPR14]. Cone [SZ22b]. Cone-Constrained [SZ22b]. Conic [BCC15]. Connecting [JT20]. Connection [NM13]. Consistency [Vig22]. Consistent [AHJ15, BZZ19, BWYY14, CCW19, EMP12, For20, RT17, vSDF19]. Constrained [BCV14, DHL15, HSX23, MW13, PVW17, SZ22b, ZRA14]. Constraint [AS18, ABY22, ABY23, BMS16, Sek13]. Constraints [AI15, ABY19, BCX19, CCC^{+}21, CVS13, JL15, JL20a, LMK13, NZ19, TA15, WX21, WXXY22]. Construction [MS14, Shi17]. Constructions [CC12, JL20b]. Consuming [BPY16]. Consumption [ABY22, ABY23, CD13, CCSW19, HZ16, Käl20, PW22, dRP22]. Consumption-Investment [PW22]. Contagion [AMS15, CK18, DMBPR19, FS21, GV16, HSZ17, JPZ19]. Contagious [SZ22a]. Contests [Cot21, NZ22]. Context [BEV14]. Contingent [AM11, FH23]. Continuity [DL11b]. Continuous [Ben11, BWYY14, DXZ10, DR18, GKR14, LPY21, MPR14, SS17]. Continuous- [LPY21]. Continuous-Time [DR18, MPR14, DXZ10]. Continuum [Lep16]. Contract [NZ19]. Contracting [CCY12, CF15]. Contracts [ANW23, AF15, BCV14, BL13, BR15, CF15, LSZ20, WYZ11]. Control [AMS15, BCV14, BDL11, CVS13, CMR19, HLLR16, Mon13b, VRT22, War13, YZZ23]. Controller [BZ14]. Controller-Stopper [BZ14]. Controlling [ACE19]. Controls [AF15]. Convergence [BCM10, BBH^{+}21, BFN22, BW11, CJM16, CMR18, FT22, HSZ17, Lev11, LHJ10]. Convertible [FH23]. Convex [AG20, Ara14, GKR14, PP10, TA15, FM11]. Copula [BGR20]. Corporate [CL21b]. Correction [DL11b, FL11, GS17, ZCLG16]. Correlated [GKMT10]. Correlators [BL21]. Cosine [ZO13]. Cost [DJ23, KP10b]. Costs [AKK13, BCDD21, Bic12, BS13b, BR15, BMS16, CD13, CMKS14, GO11a, HF10, Lep16, LMK13, PP10, RX20, DXZ10]. Counterparty [CCFK20]. Coupled [YZZ23]. Coupon [How12]. Covariance [GK22, PO10]. Cox [JC16]. Crash [SS19]. Credit [BGR20, BHH^{+}11, CL21b, CK11, EGS13, EGG10, FK21, GK16, HSZ17]. Criteria [AGZ22, Käl20, MZ10, Sch15]. Criterion [LP21]. Critical [BL16]. Cross [CFVS22, GS21, HN14, JS11]. Cross-Border [CFVS22]. Cross-Commodity [JS11]. Cryptocurrency [BHM20]. CSA [BGO21]. Cubature [CM12, FZ23]. Cumulative [KP18]. Currency [GS21]. Curvature [AL17, AFLZ17, FT15]. Curve [GS21, JO23]. Curve-Dependent [JO23]. Curves [FT15, GPSS15]. CVA [HLLR16].

D [CCL^{+}11]. Data [BAGM20]. Deal [Ara11, BND20]. Debt [BMNP17, CVS13]. Decision [BDH10]. Decomposition [FJL11]. Deep [DP20, JO23, LLM^{+}23, SZ21, TW20].

Deep-Learning [TW20]. Default [AFLZ17, CM10, CW12, CDK10, DMBPR19, DQS17, FS21, GQS20, HSZ17, JL20b, SS15]. Defaultable [BN15, LLCMA16]. Defaults [EJJ15, GKMT10]. Degeneracy [PW16]. Delayed [BD21, JSDN11]. Delayed-Exercise [JSDN11]. Delivery [Fei19, KYKLR20, PP10]. Delta [BS23, BW11, Kol15]. Densities [CFP10, GV15]. Density [DL13, EJJ15, Mon13a]. Dependence [AZ10, GV16]. Dependent [BBC^{+}20, BGM10, HKZ17, HK17b, HSZ17, JO23, KQY22, KM23, Sap20, SZ21, TW20, ZX19]. Derivative [JS11]. Derivatives [AGLG22, BK15, CK11, DMSS20, Sab23, Sap20, ZX19]. Descent [SS17]. Design [GL15, HSX15, NZ22, RL18]. Desirable [vSDF19]. Detect [JKP11]. Detection [PBH10]. Deterministic [ABY23, BF14a, JT11b]. Deviation [BZZ19, Gul18]. Deviations [AKT19, GPY13, SS15]. Difference [ARS21, BCC15]. Differences [GR12]. Different [CL21b, GMBB20]. Differential [CM21, CSS20, CM12, DB20, Hep10, SZ21, IR21]. Diffusion [BAGM20, BJ12, BS23, BL16, CM17, CLZ18, CFP10, DL11a, DL11b, Hep10, JLS20, KM23, QC19, SZ22b, ZL22, LHJ10]. Diffusions [BDG18, BWZ22, CN11, DF18, JT20]. Diffusivity [HR17]. Dimension [ACLP14, Pap13]. Dimensional [BEST23, BWZ22, BCN23, BK15, BTW11, CLP18, El 13, RBG21]. Direct [MSW21]. Directional [CFRT16]. Directions [EMW21]. Dirichlet [HR18]. Discontinuous [AN15, BCX19, DK18]. Discounting [BGO21, LLP23]. Discrete [BZZ19, BWYY14, BC22, BF14b, CKT17, CMKS14, DZ23, EGS13, Fri14, GKR14, How12, JV11, Pap13, Rás15, RTY23, RT17]. Discrete-Time [Rás15, RTY23].

Discretely [FMM11b]. Discretization [BFN22, CFRT16]. Dispersion [GPS22]. Displaced [BJ12]. Displaced-Diffusion [BJ12]. Distance [BW23]. Distribution [ABY19, KP18, vSDF21]. Distributions [NPS17, RS12]. Diversification [DV15, Qi23]. Dividend [ABY19, CLZ18, CVS13, CY21, How12]. Dividend/Coupon [How12]. Dividends [AAM22, CCY23, JV11]. Does [AG20]. Dominance [WX21, WXXY22]. Double [CASB22, CO11]. Double-Execution [CASB22]. Down [TY12]. Downgrade [CL21b]. Drawdown [AS18, ABY19, ALMY23, Sek13]. Drawdowns [CNP12]. Drift [EV16, HKMR20]. Drifts [CM17]. Driven [BGSS23, EGS13, FTT10, HX19]. Driver [GP15]. Dual [Ben11, BBC16, FM11, HLLR16, Lel18, Lev18, MT17, Mon13b, PW22, RS14, Rog10, SZH13]. Duality [BD18, CKT17, CMKS14, HH10, Kir15, TA15]. Due [GS17]. Duesenberry [Lon20]. Duopolistic [Lud11]. Dupire [ÖH23]. Duration [JT20]. Dynamic [BL15, BCC15, BND20, CF15, COW19, COW22, CCFK20, CK11, EM21, FMM11a, FS21, HW14, HZK17, JPZ19, KS10, MT17, MWZZ22, NS21, VFL23, vSDF21]. Dynamical [LLM^{+}23]. Dynamics [BDG22, CdL13, CM21, EJM21, HK17b, KM23, LLS14, ZCLG16].

Easy [RW12]. Economic [BBC^{+}20, BGH21, EM21, LN22]. Editors [CS10]. Editors-in-Chief [CS10]. Effect [FT15, HRW13]. Effects [LPY21]. Efficiency [JL15, JL20a]. Efficient [BPO18, DT13, GKMT10, Kir15, Kir16, ZO13]. Eigenportfolios [AHPP22]. Eisenberg [FPR^{+}18]. Elastic [HSX15]. Electricity [CFVS22, KYKLR20]. Element [HR18]. ELMMs [BMMB18]. Emission [HS12]. Emissions [Lud11]. Empirical [GGM17].

Endogenous [CW22b, LPY21, NZ19, Zhi23]. Endowments [BCMS22]. Energy [BCV14, BCN23, BK15, Sab23, TT18]. Enhanced [CM10]. Enlargement [CG20]. Entropy [MGH18, Qi23, Tia23]. Envelope [DHOR11]. Environment [BPR21, FH18]. Equation [CM21, IR21]. Equations [BZ20, BKX12, BCC15, BHH^{+}11, CCL^{+}11, CSS20, CM12, DB20, FZ23, HK17a, HK19, Hep10, SZ21, SSZ16]. Equilibria [SSZ17]. Equilibrium [BD18, BLD17, CW22b, EM21, FT22, KHOL10, LLLY18, LX15, LLX19, Lon20]. Equities [AHPP22]. Equity [GV16]. Equivalents [VRT22]. Ergodic [LZ17]. Ergodicity [HR17, ZRA14]. Erratum [COW22, HK19]. Error [Gas23, Gul10, RR10]. Escrow [SW22]. ESO [LW15]. Estimate [PO10]. Estimates [Gul10]. Estimation [BAGM20, FW18, MP23, RS12]. ETF [AZ10]. ETFs [AHJ15, FLGL18]. Euler [CJM16, CMR18]. European [GQS20, How12, JSDN11, MS14, ZO13]. European-Style [ZO13]. Event [GL22]. Eventual [FT15]. Evolution [BHH^{+}11, HK17a, HK19]. Exact [FW18, GKMT10, NM13, VG22]. Examples [GPSS15]. Exchange [CJJ20, CSB21, CMR18, Kar15]. Execution [AKU21, AS10, BL15, BDG22, CL23, CJ16, CASB22, GR14, GZ15, HW14, KLA20, KP10b, PSS11, Sch15]. Executive [HKMR20]. Exercise [BCV14, Ben11, HKMR20, JSDN11, JV11]. Existence [AF15, FTT10, IR21]. Exotic [TD22]. Expansion [Aly14, AFLZ17, BLP16, Lel18, Mon13b, TY12]. Expansions [FLGL18, KP16, MSW21, PPR13, ZO13]. Expectations [HO11]. Expected [BMS16, DJ23, LLX19, WX21, KP10a]. Explain [BOL14]. Explicit [CC12, CJM16]. Exponential [Ant14, FLF12, FLGL18, LLP23, RS14, ZO13, DZ23]. Extension [MS10]. Extensions [ÖH23]. Extreme [Gul10, Pap18]. Extreme-Strike [Pap18].

Factor [DL11a, DP20, FHS22, LLL21, LZ17, RX17]. Factors [BZ20, BK18]. Fast [FL11, FJL11, FH18, GS20, Guy22, TD22, FFF10, Kir15]. Fast-Varying [GS20]. Fee [JLS20]. Feedback [CSS15]. Fees [BPR21, Wan22]. Field [FT22, GPY13, GGR23]. Fill [CSB21]. Filtering [BAGM20, EMW21, PO10]. Filtering-Based [BAGM20]. Filtration [CG20]. Finance [BCC15, CD12, FJO21, GGM17, GR12, Li22]. Financial [BCM10, BBC^{+}20, BN15, BMMB19, CJM16, CK18, DMBPR19, Fei19, FH23, FL13a, FT20, Fri14, HS12, KV19, LN22, MGH18, Shi17]. Financialization [CSS15]. Financing [MS14]. Finite [BTW11, Bic12, FT22, GR12, GLZ20, HR18, LS21, MPR21, MPR14]. Finite-Dimensional [BTW11]. Fire [BF19, DMBPR22]. Firm [PVW17]. Flexible [GG18]. Flow [BL15, Jus21]. Flows [BMMB18]. Follow [GTW19]. Following [DZZ10]. Foreign [CJJ20, CSB21, CMR18]. Form [CJ16, CCL^{+}11, FK21]. Formation [ABY22, ABY23, BN15]. Formats [BEST23]. Forms [HR18]. Formula [CDK10, DF18, GS17, ÖH23, SSV18]. Formulas [CKT17, Gul10, Guy22]. Formulation [LT19]. Forward [AKK13, Ant14, AGZ22, AS22, BBK12, GK10, JR13, JR15, Käl20, LLL21, LZ17, SSZ16, dRP22]. Forwards [KYKLR20]. Fourier [Kir15, BEV14, FO11, HZ10, ZO13]. Fourier-Based [FO11]. Fractional [BD20, DB20, FH18, GS17, GNR19, GJRS18, Gul18]. Frame [Kir15]. Framework [AHO20, CKN18, DZ23, FT20, GK16, GS21, HR17, KYKLR20, LSZ20].

Free [CL21b, CD22, Fon23, GJMN16, NS21, NJZB23, SSV18, LHJ10]. Frequency [CJR14, FP15, HK18, HMKSY20, JP15, SSZ17]. Frictions [Bur16]. Full [HKMR20]. Fully [BND20, HK17b, KM23]. Fully-Dynamic [BND20]. Functional [CW22a, SSV18]. Functionally [Str14]. Functionals [BKMMS21, KS10, TD22]. Functions [BMZ11, FL13a, Gul10, MT17, Sch15, WXXY22, YZZ23]. Fund [AGZ22, NZ19]. Fundamental [DP20]. Funding [BR15, CSS20]. Future [Guy22]. Futures [AG20, BS13b, PBH10, WYZ11].

Gain [BP13]. Gain-Loss [BP13]. Galerkin [GG18, SZ21]. Game [DQS17, GGR23, LLL21, SSZ17]. Games [BDH10, Lud11, dRP22]. Gamma [BW11]. Gas [War13]. Gaussian [BGR20, DZ23, Yam22]. General [BZ20, BD18, BKMMS21, BDL11, CC16, CLZ18, CKN18, DT13, HR17, PSS11]. Generalizations [Str14]. Generalized [AN15, GJMN16, KP18, NW21, Sek13]. Generated [BFS11, RX20, Str14]. Generation [FÖ21, NJZB23]. Geometric [Qi23]. Given [CC12]. Globally [Lev11]. Glosten [LX15]. GMxB [AF15]. Goal [BPY16]. Good [Ara11, BND20]. Gradient [SS17]. Growth [QC19, Zhi23]. Guaranteed [GR14, HZK17].

Habit [ABY22, ABY23]. Habit-Formation [ABY22, ABY23]. Harmonic [De 21]. Hawkes [AJ15, BGSS23, HX19, ZRA14]. Hearts [Ara14]. Heat [AFLZ17, FZ16]. Hedge [CGJ19]. Hedging [AM11, AHO20, AKK13, BHZ15, BCR21, BL13, BR15, BS23, BBC16, BMS16, BW11, Bur16, CFRT16, CDP21, CN11, CCFK20, CL21a, CW12, CKT17, CK11, CO11, EGLO21, GS20, GS21, Kol15, PP10, RR10, Rog10, TA15]. Heston [AN15, BD20, BGM10, BHSW15, CPZ22, DB20, FO11, FFF10, FJL12, FL11, GO11b, GJRS18, JR13, JS19, LT19, MSW21, RTY23]. Heterogeneous [FS21, Lon20, MWZZ22]. Hierarchical [BEST23]. High [ACLP14, BEST23, BCN23, CJR14, DM22, El 13, FP15, HK18, HMKSY20, JLS20, JP15, Li22, RBG21, SSZ17]. High-Dimensional [BEST23, BCN23, El 13, RBG21]. High-Frequency [HK18, HMKSY20, SSZ17]. High-Order [Li22]. High-Watermark [JLS20]. Higher [BW11]. Hilbert [Hep10]. HJB [SSZ16]. HJM [DT13]. Homogeneous [CN11]. Homothetic [LLL21, LZ17]. Honest [EE22]. Horizon [GLZ20, LLLY18, LZ17]. Horizons [KN18, LS21]. Hybrid [CMR18, CMR19]. Hyperbolic [FZ16, KP18]. Hypergeometric [CMP21].

Ill [SSZ16]. Ill-Posed [SSZ16]. Imbalance [BL15]. Impact [BV19, DM22, FJS22, GZ15, KP16, NV22, NPS17, RX20, SSZ17]. Imperfect [CCY12, DQS17, HMKSY20]. Implied [ARS21, BLP16, CMN17, CC12, CDK10, DHJ17, De 21, EJM21, FJL12, Gul10, GV15, JR15, JKR18, LS16, LLCMA16, NPS17, NJZB23, Teh16]. Improving [CSB21]. Impulse [BDL11]. Inclusion [LDDD21]. Incoming [BF14b]. Incomplete [AM11, BDH10, GQS20, NZ13, SSZ16, Tia23]. Inconsistency [HW21]. Inconsistent [BWZ22, CLZ18, Sch15]. Independent [Bur16, ZŽ10]. Indices [GV16]. Indifference [BD21, BZ14, BND20, DM22, VG22]. Indirect [Mos21]. Individual [FPR^{+}18]. Induced [Ara11, BMMB18]. Induction [Lev18]. Inferring [Cot21]. Infinite [BK15, LZ17]. Infinite-Dimensional [BK15]. Inflation [DMSS20]. Inflation-Linked [DMSS20]. Information [AHO20, ACE19, AMS15, BD21, BUV12, GL22, HW14, HKMR20, LSWY18, Pap13]. Informational [JL15, JL20a]. Ingersoll [JC16]. Inhomogeneous [DMBPR19, EGS13]. Initial [BGO21]. Initiation [HZK17]. Injections [NZ19]. Insider [CD12, EE22]. Insiders [CD22]. Instruments [DVW11, HS12]. Insurance [BPY16, BS13a]. Integrable [SS19]. Integral [BCV14, FZ23]. Integrated [BGH21]. Integration [AN15, GGM17]. Integro [Hep10, IR21]. Integro-Differential [Hep10, IR21]. Intensities [CDK10]. Interbank [AMS15, FPR^{+}18, FS21, FI13, KV19]. Interbanking [BC15]. Interest [BGSS23, BHM20, GG14, GO11b, ÖH23, ZX19]. Interpolation [BJ12, GKS20]. Intertemporal [Vig22]. Intractable [HX16]. Intraday [BAGM20, CSB21]. Invariant [BKMMS21]. Inventories [CL23]. Inventory [HMKSY20]. Inverse [Käl20]. Inversion [CDK10, MSW21, AG20]. Inverting [FL13a]. Investment [ABY22, BV19, BK22, BCMS22, BHY19, Bic12, CD13, CCSW19, CVS13, GLZ20, JLS20, Käl20, LLLY18, LLL21, MT17, NZ13, PW22, PVW17, Rás15, RX17, Sek13, VG22, dRP22, vSDF21]. Investment-Consumption [dRP22]. Investor [ACE19, BUV12, DV15]. Investors [GTW19]. Iteration [RW12].

Job [LW15]. Joint [GLOW22, Guy22]. Jump [AFLZ17, BK22, BAGM20, BS23, BL16, CM10, DL11a, DL11b, GK16, Hep10, JLS20, KM23, NPS17, LHJ10]. Jump-Diffusion [BAGM20, BS23, DL11a, DL11b, Hep10]. Jump-to-Default [AFLZ17]. Jumps [BZ14, BCN23, LW15, LPY21].

Kelly [LP21]. Kernel [AFLZ17, FZ16]. Kind [Lep16]. Knightian [BD18]. Knowledge [QC19]. Kusuoka [Yam22]. Kyle [GMBB20, KHOL10].

Lag [HK18]. Lagrangian [War13]. Large [AKT19, ACE19, BGR20, GPY13, Gul18, HK17a, HK19, HK17b, JL18, SS15]. Large-Tick [JL18]. Last [CNP12]. Latency [CSB21]. Law [BKMMS21, HK17b]. Law-Invariant [BKMMS21]. Lead [CF15, HK18, LDDD21]. Lead-Lag [HK18]. Learning [CCC^{+}21, CL21a, JPWT22, MP23, Pun21, TW20]. Lending [FI13]. Level [DK18, FT15, JL18, KP10a]. Level-I [JL18]. Level-Slope-Curvature [FT15]. Leverage [LLM^{+}23]. Leveraged [AHJ15, AZ10, FLGL18]. Lévy [BEV14, BPO18, CMN17, EGS13, FL13b, FLF12, FLGL18, GG18, JL13, JS11, LLP23, Lev18, LP21, PPR13, RS14, ZO13]. Lévy-Based [JS11]. Lévy-Type [JL13]. Liabilities [BS13a, FPR^{+}18]. LIBOR [BJ12, GPSS15]. Life [BPY16, BS13a]. Lifelong [HZK17]. Lifts [DJ23]. Limit [AJ15, AS10, CGJ19, CdL13, CM21, FT22, GLFT12, HN14, HK17b, HX19, JL18, KM23, PSS11, SSZ17, SV17, ZCLG16, ZRA14]. Limit-Order [PSS11]. Linear [CM17, DM22, FJO21]. Linearization [DB20]. Linked [DMSS20]. Linking [DHL15]. Lipschitz [CJM16]. Liquidation [EV16, GLFT12, JL18, KP10b]. Liquidity [AKU21, Alm12, AKK13, BF14a, BMMB18, JP15, LLP11]. Loans [GHK21]. Local [AM11, AG20, AFLZ17, BFS11, BPO18, CM10, CCC^{+}21, CMR18, CMR19, CKN18, DF18, FLGL18, JKR18, KP16, LS16, ÖH23, PPR13, PZ16]. Local-Stochastic [AFLZ17, CMR19, LS16]. Locally [BL13]. Log [BHP21]. Log-Modulated [BHP21]. Long [AJ15, AKT19, KYKLR20, RX17, Sek13]. Long-Term [KYKLR20, RX17, Sek13]. Long-Time [AJ15, AKT19]. Lookback [GL22]. Looping [JPZ19]. Loss [BP13, BMS16]. Low [CJR14, GKS20, RBG21]. Low-Dimensional [RBG21]. Low-Rank [GKS20]. LSMC [FJJ20]. LT [ACN13]. Lunches [CD22].

Machine [CL21a]. Magic [GGM17]. Make [BPR21]. Make-Take [BPR21]. Maker [BPR21]. Makers [BDH10]. Making [Jus21]. Malliavin [ATL12, AN15, Mon13b, Sap20, TY12]. Managed [Pun21]. Management [AGZ22, BMNP17, DL11a, EMP12, HMKSY20]. Manager [NZ19]. Managing [DMBPR19, WYZ11]. Manipulation [ASS12]. Manipulations [AS10]. Many [dRP22]. Maps [FM11]. Margin [BGO21, Fei22]. Market [ACE19, BPR21, BJ12, BFZ21, CGJ19, CT15, CdL13, DQS17, Fri14, GQS20, Jus21, LS21, LLCMA16, PBH10, SSZ17, ZCLG16, Zhi23, dRP22, KP10a]. Market-Adjusted [BFZ21]. Market-Maker [BPR21]. Markets [AM11, BCV14, BCN23, BDH10, BEV14, BOL14, BK15, BN15, Bur16, CJQ16, CSB21, CSS15, CMR18, FHS22, HKZ17, HS12, Lud11, NZ13, Rás15, SZ22a, SSZ16, SV17, Tia23, WX21]. Markov [CLL21, DR18, FP15, MPR14, ZL22]. Markovian [CdL13, HR17, SV17, ZCLG16]. Markowitz [Arm18, DXZ10, El 13, HX16, JMP21]. Martingale [DHL15, JKR18, LR16, SS19]. Martingales [CC12, SZH13]. Mass [DHJ17]. Master [SSV18]. Mathematical [Li22]. Matrix [GK22, PO10, RX17]. Maturities [KV19]. Maturity [BL16, CC16, FLF12, FMM11b, JR13, PZ16, ZŽ10, FFF10]. Maturity-Independent [ZŽ10]. Max [KHOL10]. Maximization [AD23, BCDD21, BMZ11, BCX19, BS13b, CL21a, DZ23, FMM11a, LLX19, MW13, VF16, WX21, Zho21]. Maximize [AG20]. Maximum [FMM11a, MGH18, Qi23]. Maxmin [LLX19]. May [VFL23, LDDD21]. Mean [ALMY23, BZZ19, BHY19, CM17, CCW19, De 21, DVW11, For20, FL11, FJL11, FH18, FT22, GPY13, GGR23, HSX15, HX16, HSX23, LLLY18, Liu15, MS14, Sch15, SZ22a, SZ22b, vSDF19, vSDF21, FFF10]. Mean-Quadratic [vSDF19]. Mean-Reverting [CM17, DVW11, FL11, FJL11, FH18, Liu15, FFF10]. Mean-Self-Financing [MS14]. Mean-Standard [BZZ19]. Mean-Variance [ALMY23, BHY19, CCW19, HSX15, HX16, HSX23, LLLY18, Sch15, SZ22a, SZ22b, vSDF19, vSDF21]. Measure [BOL14, BP13, WBT15]. Measurement [RL18]. Measures [Ara14, BFZ21, CG20, DR18, DF21, FRW17, FTT10, HH10, ZŽ10]. Merton [BK22, BST10, PW16]. Message [CS10]. Metamodel [BGR20]. Method [ACN13, ACLP14, CMR19, CM12, FO11, FZ23, Kir16, Li22, Mon13b, SZ21, ZX19, HZ10]. Methods [ATL12, BCM10, BDN15, BHU18, BEV14, BHSW15, FJJ20, HK18, HR18]. Microcredit [LDDD21]. Microstructural [RR10]. Milgrom [LX15]. Minimal [WXXY22]. Minimalist [FLW23]. Minimization [AM11, BS13a, Fri14]. Minimize [ALMY23]. Minimizing [KP10a]. Minimum [BFS11]. Misspecified [JT11a]. Mixed [GV15]. Mixing [FJJ20]. Mixture [AS22, MPR21]. Model [AKU21, AAM22, AKT19, AN15, ANW23, BHZ15, BGM10, BJ12, BGH21, BS23, BGR20, BL16, BMNP17, Bur16, CJQ16, CDJ17, CCY23, CSS15, CLP18, CPZ22, CMP21, CDK10, CM21, CMR18, CMR19, DZZ10, DB20, DL11a, DJ23, FO11, FP15, FJL12, FZ16, FL11, FI13, GMBB20, GPY13, GG14, GGR23, GQS20, GO11b, GJRS18, HK17b, HR18, HX16, JS19, JLS20, JSDN11, JC16, KP16, KM23, LT19, LMS22, LS21, Liu15, MSW21, NS21, QC19, RTY23, SSV18, VF16, DXZ10, FFF10]. Model-Free [NS21, SSV18]. Model-Independent [Bur16]. Modeling [BBK12, CCC^{+}21, EJJ15, GV16, GLOW22, RL18, SV17]. Modelling [BHH^{+}11, ZRA14]. Models [AE19, AD23, AS10, AL17, Aly14, BLP16, BCM10, BD20, BHP21, BFN22, BKX12, BAGM20, BYY12, BGSS23, BPO18, BHSW15, CM17, CMN17, CFR13, CT15, CW12, CGSF23, CKN18, DMSS20, DL11b, DP20, DT13, EGS13, EFGR19, FL13b, FLF12, FLGL18, FTT10, Fon23, FZ17, FJL11, FN18, Fri14, FK21, GG18, GS12, GPSS15, GV15, Gul18, Guy22, HK17a, HLLR16, Hep10, HJT20, HR17, JMP21, JL13, JKR18, JS11, JT11a, JT11b, KHOL10, LLP23, Lev18, LLL21, LZ17, LPY21, MPR21, MKPS12, NW21, PPR13, PZ16, RBG21, RT17, RX17, Sap20, SZ22b, Shi17, ZL22, HK19]. Modulated [BHP21]. Monetary [KS10]. Money [EFGR19]. Monitored [FMM11b]. Monotone [HSX23, MZ10, SZ22b]. Monte [AN15, BHU18, GR12, MPR14, ZX19]. Mortality [BBK12]. Mortgage [JC16]. Motion [GNR19]. Moving [BTW11]. Multi [BPR21, BDG22]. Multi-asset [BDG22]. Multiasset [AKT19, CD13]. Multicurrency [GG14]. Multidimensional [JLS20]. Multientrant [Cot21]. Multifactor [AE19, GS12, KYKLR20]. Multilayered [Fei19]. Multilevel [AN15, BDN15, GR12, ZX19]. Multiperiod [BW23, For20]. Multiple [AI15, ACLP14, ANW23, Ben11, BC22, BK18, EGLO21, GS21, GPSS15, GK22, KV19]. Multiplicative [GZ15]. Multiscale [BCM10, FHS22, Sap20]. Multivariate [AD23, ACDP18, GMBB20, JMP21, MS10, MKPS12].

Nash [SSZ17]. Natural [War13]. Near [BL16, Lev11]. Nested [BHU18]. Net [HSX15]. Network [BCN23, CK18, Fei19, SZ21]. Networks [AFM20, BBC^{+}20, BMMB19, BC15, DMBPR19, KV19, LLM^{+}23]. Neural [BCN23, LLM^{+}23, SZ21]. Neutral [HS12, Mon13a]. No [BND20, Lep16, MM22]. No-Good-Deal [BND20]. Node [AFM20]. Noe [FPR^{+}18]. Noise [CW22b]. Non [CJM16]. Non-Lipschitz [CJM16]. Nonconcave [GLZ20, Rás15]. Nonlinear [CM17, CF15, CM12, GQS20, NM13]. Nonparametric [ATL12]. Nonquadratic [AM11]. Nonsmooth [ABRR21, BMZ11, GP15, GLZ20, HRW13, ZL22]. Nonstationary [MGH18]. Nonuniformly [SS19]. Normal [CFR13, Sab23]. Note [BD21, BCDD21, ZCLG16]. Numbers [HK17b]. Numerical [ACLP14, BLP16, BCR21, CCL^{+}11, Gas23, Li22, PVW17, PBH10, VFL23]. Numerics [EGLO21].

Obligations [Fei19, FH23]. Observable [EM21]. Observation [LPY21]. Observations [CCY12]. Obstacle [ABRR21]. ODEs [Yam22]. One [BWZ22, CLP18, FLW23, PSS11, Zha23]. One-Dimensional [BWZ22, CLP18]. One-Sided [PSS11]. Opportunities [JT11a, PVW17]. Optimal [AKU21, ACLP14, AAM22, AS10, Alm12, ANW23, ABY19, ABY22, ABY23, Ant14, ALMY23, AF15, BPR21, BF14a, BV19, BK22, BCV14, BBH^{+}21, BDG18, BCDD21, BCN23, BL15, BCMS22, BDG22, Bic12, BDL11, BMNP17, CFRT16, CDP21, CL23, CFVS22, CCY23, CT15, CD13, CLZ18, CCSW19, CVS13, CMP21, CY21, DHL15, DVW11, DV15, EV16, FH18, FJS22, GS20, GL15, GLZ20, GLFT12, GZ15, GLOW22, HZ16, HZK17, JL18, JLS20, JC16, Jus21, KLA20, KP10b, KS10, KN18, LLP23, LLP11, LL11, LHJ10, LS21, LPY21, MWZZ22, NZ13, NZ19, NV22, PSS11, PBH10, Rás15, RX17, SZH13, Sch15, Sek13, TT18, VG22, War13, vSDF21]. Optimality [Vig22]. Optimization [ACE19, BW23, BD20, BHY19, BF19, CW22a, FHS22, JPZ19, LS16, MP23, Pap13, Vig22, vSDF19]. Optimized [VRT22]. Option [ACN13, AKT19, Aly14, BFS11, CGJ19, CLP18, CCL^{+}11, CLP11, DL13, FJL11, FR14, GGM17, GG18, GKS20, GO11a, HKMR20, Hep10, HF10, IR21, JT20, Kir15, Kir16, LT19, MKPS12, NS21, RS12, HZ10]. Options [ATL12, ABRR21, AHJ15, BQY22, BEST23, BZ14, BHZ15, BDN15, Ben11, BDH10, BYY12, BL13, BTW11, BK18, BBC16, BHSW15, CPZ22, CO11, DHL15, DK18, DL11b, DQS17, EGLO21, FO11, FJJ20, FL13b, FLGL18, FZ16, FMM11b, GL22, GKR14, GQS20, HJT20, How12, HRW13, JSDN11, JV11, Kar15, Kol15, Lel18, LL11, Lev18, Liu15, MS14, PPS18, Pap18, PZ16, RW12, Rog10, TD22, ZO13, Zho21, LHJ10]. Order [AJ15, AKU21, AS10, ASS12, BF14a, BL15, BF14b, BW11, CdL13, CM21, GKR14, HN14, HK17b, HX19, HR17, JL18, Jus21, KM23, Li22, PSS11, Shi17, SV17, ZCLG16]. Ordering [AG20]. Orders [CGJ19, GLFT12, LLP11]. Orlicz [Ara14, LN22]. Ornstein [BDG22]. Orthogonalization [LR16].

Pairs [CCW19]. Pairs-Trading [CCW19]. Parabolic [CCL^{+}11]. Parameter [DV15]. Parameters [AKU21]. Parametric [DB20, GGM17, GKS20]. Parametrix [CFP10]. Parareal [PPS18]. Parisian [DL13]. Parities [Kar15]. Partial [ACE19, AMS15, BUV12, CM21, HKMR20, Hep10, LSWY18, Pap13, SZ21, TA15]. Particle [CMR19]. Parts [AN15]. Passage [CNP12]. Path [AZ10, CASB22, DJ23, FÖ21, HSZ17, KLA20, KQY22, QC19, Sap20, SZ21, ZX19]. Path-Dependence [AZ10]. Path-Dependent [HSZ17, KQY22, Sap20, SZ21, ZX19].

Patterns [BF14a]. Payments [How12]. Payoffs [AN15, CGJ19, HRW13]. PDE [FJJ20, MT17, NM13]. PDEs [FJO21, JO23]. Penalized [HSX15]. Penalty [CLZ18, HRW13]. Performance [AGZ22, AS22, BP13, BLD17, LLL21, LZ17, MZ10, SSZ16, Wan22]. Performances [Ant14]. Period [KYKLR20]. Periodic [LPY21]. Periodic-Observation [LPY21]. Perpetual [CDP21, GL22]. Perpetuals [ACEL23]. Persistent [Jus21]. Perspective [CJR14]. Perspectives [LN22]. Phase [PBH10]. Physical [Fei19, PP10]. Players [dRP22]. Point [EGG10, GK22, HKMR20]. Points [GGM17, KQY22]. Poisson [FTT10]. Policies [BYY12, TT18]. Policy [RW12]. Polynomial [BL21, FÖ21, KYKLR20]. Pools [LLP11, SS15]. Poor [BP13]. Portfolio [AI15, AS18, ASS12, ACE19, BD20, BWYY14, BGZ22, BGR20, BHH^{+}11, CW22a, CMP21, CK11, EMP12, EGG10, FH18, FHS22, GLFT12, HW14, HSX15, HX16, JMP21, JPZ19, KP10b, KN18, LS21, LMK13, LS16, MZ10, Pap13, PO10, PW16, Pun21, PBH10, Qi23, RL18, SSV18, SZ22a, SZ22b, TW20, vSDF19]. Portfolios [BCDD21, DVW11, EGS13, El 13, HK17a, HK19, MP23, RX20, Str14]. Posed [SSZ16]. Posedness [NW21]. Positions [AKK13, CD22]. Positive [ASS12, DHJ17, LPY21]. Positivity [FTT10, FT15]. Possibility [CW12]. Power [AS22, BEV14, BOL14]. Predictable [GKR14]. Preferences [CLZ18, Vig22]. Premium [ALMY23, BOL14, GV16]. Prepayment [JC16]. Presence [DV15]. Price [AG20, AS10, ASS12, BV19, BL16, CJ16, CSB21, CdL13, DM22, FJJ20, FJS22, GGR23, GV15, GZ15, KP16, KM23, LW15, Lev11, NS21, NV22, RBG21, SSZ17, ZCLG16]. Prices [BC22, CMKS14, Fei22, FJL11, FR14, GTW19, JT20, LT19, RS12, VG22, ZRA14, Zhi23]. Pricing [ACN13, AHJ15, AHO20, AKT19, Aly14, Ant14, BD21, BQY22, BEST23, BZ14, BDN15, Ben11, BOL14, BK15, BL13, BTW11, BLD17, BND20, BHSW15, CLL21, CLP18, CL21b, CCL^{+}11, CW12, CKT17, CLP11, CM12, DL13, DM22, EGLO21, FL13b, Fon23, FMM11b, GGM17, GG18, GS12, GKS20, Gul10, Hep10, HF10, HS12, HO11, HSZ17, IR21, Kir15, Kir16, Lel18, Lev18, MPR14, MKPS12, RW12, Sab23, Sap20, Tia23, TD22, ZO13, HZ10]. Primal [Ben11, MT17]. Primer [ACEL23]. Principal [AHPP22, BH16, MR18]. Principal-Agent [BH16, MR18]. Principle [ALMY23, FMM11a, Gul18, MGH18, Tia23]. Principles [MW20]. Priors [BC22]. Probabilistic [ACLP14]. Probability [ALMY23, Cot21]. Problem [ABRR21, ASS12, BH16, BK22, BCV14, BZZ19, BST10, BCR21, COW19, COW22, CL21b, CVS13, CY21, DHL15, GLZ20, HSX23, Käl20, LLLY18, MR18, MW13, NZ13, PW16, TA15]. Problems [ACLP14, BW23, BZ14, BCN23, BMZ11, GL15, HLLR16, LR16, MT17, Mon13b, Vig22, War13, YZZ23]. Process [AJ15, CLZ18, EM21, FL13b, MPR14, Mos21]. Process-Based [AJ15]. Processes [Ara14, AS22, BEV14, BL21, BGSS23, CLL21, CNP12, EGS13, EGG10, FTT10, HX19, LLL21, LZ17, LP21, RS14, Sab23, SSZ16, ZO13, ZRA14]. Producer [TT18]. Products [SZH13]. Programming [VFL23]. Progressive [CG20]. Prohibition [BWYY14]. Projection [TD22]. Propagation [FK21]. Properties [AGLG22, Liu15]. Proportional [BCDD21, Lep16, RX20]. Prospect [KP18]. Purchase [LL11]. Purchasing [BPY16]. Push [TY12]. Push-Down [TY12]. Put [BL16, CDP21, JV11, MS10]. Put-Call [MS10].

Quadratic [CL23, CFR13, GS12, GKR14, JMP21, KP16, Qi23, vSDF19]. Quadrature [AN15]. Quantifying [AHO20]. Quantile [BCR21, BBC16, WXXY22]. Quantiles [BP22, FLW23]. Quantum [FJO21]. Quasi [DB20, FM11, RS14]. Quasi-convex [FM11]. Quasi-Linearization [DB20].

Radner [CW22b]. Random [BKMMS21, BF14b, GGR23, HSX23, LLLY18, LP21, Yam22]. Random-Supply [GGR23]. Randomization [FMM11b, Lev11]. Randomized [BBH^{+}21, JS19]. Range [LSWY18]. Rank [GKS20, HKZ17]. Rank-Dependent [HKZ17]. Rao [Qi23]. Ratcheting [AAM22, ABY19]. Rate [BFN22, CJM16, GK10, ZX19, LHJ10]. Rates [ABY19, BGSS23, BHM20, CMR18, CMR19, Fon23, Gas23, GG14, GO11b, ÖH23]. Rating [CL21b]. Ratio [AS18, BP13, LS16]. Rational [DMSS20]. Ratios [CSB21]. Reach [BPY16, KP10a]. Real [ABRR21, BDH10, BYY12]. Realization [KQY22]. Realized [El 13]. Recalibration [RT17]. Recombining [BDG18]. Recover [EM21]. Recovery [Mon13a]. Recursive [DL13, FMM11a]. Reduce [GO11a]. Reduced [BHSW15, CLP11]. Reduction [ATL12, BHU18, MPR14, Pap13]. Reference [KQY22]. Regime [BYY12, CCW19, DZZ10, MS14]. Regime-Switching [BYY12, CCW19, MS14]. Regression [BHU18]. Regression-Based [BHU18]. Regularity [JV11, MT17]. Regularization [RS12]. Regularized [MP23]. Regulatory [MW20]. Reinforcement [JPWT22]. Reinsurance [ALMY23, CCFK20, MWZZ22]. Related [MT17]. Relative [AGZ22, BLD17, Cot21, Pun21, Tia23, dRP22]. Relative-Volatility-Managed [Pun21]. Relaxations [HW14]. Relaxed [BUV12].

Renewal [FP15]. Repayment [GHK21]. Replicating [CGJ19]. Replication [BC22]. Representation [BBC16, De 21, FM11, LZ17, Sap20]. Representations [CLP18, KS10, MT17]. Resilience [ASS12]. Resolution [PW16]. Respect [BW23]. Response [EMW21]. Return [DJ23, PW22, Qi23]. Returns [AZ10, TW20]. Reverse [LMS22]. Reverting [CM17, DVW11, FL11, FJL11, FH18, Liu15, FFF10]. Reward [NZ22]. Right [VFL23]. Risk [AI15, AM11, AAM22, AFM20, AKK13, Ant14, Ara11, Ara14, ACDP18, BBC^{+}20, BOL14, BS13a, BF19, BND20, BC15, BGR20, BFZ21, CG20, CCFK20, DL11a, DR18, DF21, DM22, DV15, El 13, EGG10, FRW17, FKV12, FP15, Fon23, For20, Fri14, FK21, GL22, GPY13, GK16, GV16, HH10, HKZ17, HZ16, HS12, HSZ17, JPWT22, JPZ19, KP10b, LR16, LSZ20, LW15, LSWY18, MW20, Mon13a, NZ22, Sch15, WBT15, WYZ11, ZŽ10]. Risk-Aware [JPWT22]. Risk-Dependent [BBC^{+}20]. Risk-Free [Fon23]. Risk-Indifference [BND20]. Risk-Minimization [AM11, BS13a]. Risk-Neutral [HS12, Mon13a]. Risk-Sensitive [DL11a]. Risk-Sharing [LSZ20]. Risk-Taking [NZ22]. Risky [EGS13]. Robust [AHO20, BLP16, BGZ22, BFZ21, COW19, COW22, CL21a, CKT17, CO11, EGLO21, HX16, JPWT22, Lep16, LLL21, PW22, VF16, YZZ23]. Robustness [BS23, DHOR11, LN22]. Role [CD22, GV16]. Ross [JC16]. Rough [AE19, BD20, BHP21, BQY22, BFN22, DB20, DJ23, FZ17, Gas23, HW21, HJT20, JO23, KLA20, RTY23]. Ruin [CLZ18]. Rule [FLW23].

S&P [Guy22]. SABR

[CKN18, FZ16, HR18]. Saddlepoint [HO11]. Sale [HZ16, JL15]. Sales [BF19, DMBPR22]. Sample [GK22].

Sampling [ACN13]. Scaling [HX19, ZRA14]. Scheduling [BF14a]. Scheme [BCR21, CJM16, CMR18, GG18, Shi17, LHJ10, NZ13]. Schemes [BJ12, BW11, DT13, Gas23]. Scholes [BHSW15, GS17, Teh16]. SDEs [CJM16, HSZ17, NM13]. Second [Lep16]. Selection [AI15, BWYY14, HX16, JMP21, LMK13, PO10, Pun21, SZ22a, SZ22b, TW20]. Self [MS14, RS14]. Self-Dual [RS14]. Sell [CJR14]. Selling [BWYY14]. Semi [SV17, War13]. Semi-Lagrangian [War13]. Semi-Markovian [SV17]. Semistationary [BEV14]. Sensitive [DL11a]. Sensitivities [Lev11]. Sensitivity [BW23, BGH21, FPR^{+}18]. Sentiment [ACE19]. Separability [LN22]. Sequential [GL15, MP23, RL18]. Serially [TW20]. Series [LS16, MSW21, MGH18]. Set [HH10]. Set-Valued [HH10]. Setting [BWYY14]. Several [BDH10]. Shadow [CSB21, CMKS14]. Shape [CCC^{+}21]. Shapes [DHJ17]. Sharing [Ant14, LSZ20]. Sharpe [AS18, LS16]. Short [AG20, ACEL23, BD21, BFN22, BCDD21, BWZ22, BP22, BWYY14, BC22, BGH21, BFZ21, CCC^{+}21, CD22, DP20, DM22, DZ23, EJM21, EFGR19, EE22, FS21, Fei22, FFF10, FLGL18, Fon23, FJO21, GHK21, JL15, PZ16, Sap20, SZ22b, TD22, WXXY22, WYZ11, Yam22, Zha23]. Short-maturity [FFF10]. Short-Selling [BWYY14]. Short-Term [EFGR19, WYZ11]. Short-Time [FLGL18]. Shortfall [Ara11, ACDP18, Fri14]. Shorting [Zho21]. Should [GTW19]. Shrinkage [GK22]. Sided [HN14, PSS11]. Sigma [CNP12]. Signal [NV22]. Signal-Adaptive [NV22]. Signals [BK22, FJS22]. Signature [CGSF23]. Signature-Based [CGSF23]. Signatures [CASB22, KLA20]. Simple [LP21]. Simulation [CMN17, DT13, GKMT10, RTY23]. Single [CDP21]. Skew [AGLG22, Guy22]. Skewed [KP18]. Slope [FT15]. Small [AFLZ17, BMS16, DM22, FLF12, FP15, FJL12, FZ16, JR13, KN18, LS21, LMK13]. Small-Maturity [FLF12, JR13]. Small-Time [AFLZ17, FJL12, FZ16, LS21]. Smile [AL17, AGLG22, CC16, FLF12, FJL12, JL13, JR13]. Smiles [CC12, EJM21]. Smooth [BMZ11, FW18]. Snell [DHOR11]. Solution [DL13, NZ13, TW20]. Solutions [IR21, KP16, MT17, QC19, VG22]. Solvable [CM17, NM13]. Solve [SZ21]. Solving [CM12, DB20, Yam22]. Sophisticated [Zha23]. Source [CSS20]. Space [Hep10, Lev18, MZ10]. Space-Time [MZ10]. Space-Valued [Hep10]. Spaces [BKMMS21, LN22]. Sparse [Pun21]. Spatial [RL18]. SPDEs [BQY22, GR12]. Spectral [AI15, CLP18, DB20, FJL11, Mon13a]. Speculate [CGJ19]. Split [LLP11]. Splitting [DT13]. Spread [HN14, HZ10]. Spreads [CDK10]. SPX [GLOW22, Pap18]. Squeeze [Fei22]. SSVI [EJM21]. Stability [BCDD21, BWZ22, FH23, FI13, KP16, MW13, Mos21, NW21]. Stable [Sab23]. Standard [BZZ19, GL22]. State [HK17b, KM23, MPR14]. Static [CN11, TA15]. Statistical [BDG22, Str14]. Stepping [War13]. Sticky [BGZ22]. Stochastic [AKU21, AS18, AKT19, Alm12, AL17, Aly14, AFLZ17, BCM10, BCV14, BHP21, BKX12, BCMS22, BCC15, BK18, BMNP17, BHH^{+}11, CW22a, CCSW19, CMP21, CM21, CMR18, CMR19, CSS20, CM12, CKN18, EFGR19, FZ23, FZ17, FL11, FJL11, FH18, FN18, FJS22, FHS22, GS17, GS20, GR12, GG14, GO11b, GV15, Gul18, HK17a, HK19, HLLR16, JT11a, JT11b, LLP11, LLL21, LZ17, LS21, LS16, Lud11, MT17, MKPS12, NM13, NW21, ÖH23, Sap20, SS17, VRT22, WX21, Wan22, WXXY22, War13, FFF10, HK19]. Stochastic-Local [CMR18]. Stock [GV15, HKMR20, LW15]. Stocks [JL18, LLS14]. Stopper [BZ14]. Stopping [BBH^{+}21, BDG18, BZZ19, BWZ22, DL13, GL15, KS10, LLP23]. Stops [IR14]. Storage [HF10, War13]. Strategies [BDG22, BF14b, CFRT16, CASB22, HZ16, LLLY18, LLX19, MS14, VG22, Zhi23, vSDF21]. Strategy [CJ16, CD13, CLZ18]. Strict [JKR18]. Strike [ARS21, Pap18]. Strikes [Gul10]. Strong [CJM16, FT22]. Structural [FK21, GK16, HS12, Pap18]. Structural-Form [FK21]. Structure [BGSS23, CLP18, FTT10, FW18, FJL12, JT20, LPY21, RT17]. Student [GHK21]. Style [ZO13]. Sub [FHS22]. Sub- [FHS22]. Subject [WXXY22]. Suboptimality [Kol15]. Subsistence [CCSW19]. Successive [EJJ15]. Suffocating [DMBPR22]. Super [BC22]. Super-Replication [BC22]. Superadditive [WBT15]. Superreplication [COW19, COW22]. Supersolution [FHS22]. Suppliers [JP15]. Supply [GGR23]. Surface [BBK12, NJZB23]. Surfaces [CMN17, GJMN16]. Surrogate [CCC^{+}21]. Survival [Zhi23]. Suspensions [FT20]. SVI [GJMN16, MM22]. Swap [ARS21, HSZ17]. Swaps [CLL21, LLCMA16]. Swing [BCV14]. Switching [ACLP14, BCN23, BYY12, CCW19, DZZ10, Lud11, MS14]. Symmetric [EJM21]. Symmetry [MS10]. Systematically [RX20]. Systemic [AFM20, ACDP18, BF19, BC15, BFZ21, DF21, FRW17, GPY13]. Systems [CK18, FS21, KP16, LLM^{+}23].

Tail [RL18, Vig22]. Take [BPR21]. Taking [HKZ17, NZ22]. Tangent [CMN17]. Target [CJ16]. Taxes [BST10]. Taylor [LS16]. Tempered [Sab23]. Temporary [NV22]. Tenor [EGS13]. Tensor [BEST23, GKS20]. Term [BPY16, BGSS23, CLP18, EFGR19, FTT10, FW18, FJL12, JT20, KYKLR20, RT17, RX17, Sek13, WYZ11]. Term-Structure [FW18]. Terminal [vSDF21]. Termination [LW15]. Their [BBH^{+}21, BZ14, CD22, MT17, SZH13, SSZ16, LHJ10]. Them [FLW23]. Theoretical [LLL21]. Theory [BD18, BHM20, CW22a, CGSF23, GPSS15, KP18, RT17, SSV18]. Third [Shi17]. Third-Order [Shi17]. Threshold [BYY12, GK16]. Threshold-Type [BYY12]. Thresholds [CL21b]. Tick [JL18]. Tikhonov [PO10]. Time [AJ15, AKT19, AFLZ17, BZZ19, BWZ22, BCMS22, Ben11, BGM10, BWYY14, Bic12, BC22, CN11, CLL21, CLZ18, CCW19, CKT17, CMKS14, DL13, DR18, DZ23, EGS13, EMP12, FLGL18, FJL12, FZ16, For20, Fri14, GLZ20, HW21, KN18, LS21, LLCMA16, MGH18, MPR14, MZ10, Pap13, Rás15, RTY23, RT17, Sch15, SS17, War13, vSDF19, DXZ10, KP10a]. Time-Changed [CLL21]. Time-Changes [LLCMA16]. Time-Coherent [DR18]. Time-Consistent [BWYY14, CCW19, EMP12, vSDF19]. Time-Homogeneous [CN11]. Time-Inconsistency [HW21]. Time-Inconsistent [BWZ22, CLZ18, Sch15]. Time-Inhomogeneous [EGS13]. Time-Varying [BCMS22]. Times [CNP12, JL20b]. Timing [LL11]. Total [FT15]. Touch [CO11]. Trackers [CW22b]. Tractable [CFR13, MGH18]. Trade [AKU21, AS10, CDP21, JT20, Sch15]. Trading [Alm12, BCX19, BS13b, BDL11, CD12, CJR14, CDJ17, CJJ20, CFVS22, CCW19, DZZ10, EE22, FP15, FJS22, FT20, GNR19, HMKSY20, IR14, JP15, JL20a, NZ19, NS21, NV22, TT18, HN14]. Transaction [BCDD21, Bic12, BS13b, BMS16, CD13, CMKS14, DJ23, GO11a, Lep16, LMK13, PP10, RX20, DXZ10]. Transform [GS12, Kir16, HZ10, Kir15]. Transient [BV19, NV22, SSZ17].

Transition [CFP10]. Transport [DHL15, GLOW22]. Tree [BDG18, LHJ10]. Trend [DZZ10]. Triplets [CJJ20]. Tsallis [Tia23]. Two [BS13b, DK18, HN14, NM13, NW21]. Two-Level [DK18]. Two-Sided [HN14]. Type [BYY12, Gul18, JL13, MKPS12].

U.S. [AHPP22]. Uhlenbeck [BDG22]. Uncertain [FR14, FN18, HLLR16]. Uncertainty [BHZ15, BD18, CJQ16, CDJ17, CCY23, DV15, EV16, LMS22]. Underlying [MPR14]. Understanding [AGLG22]. Unified [BGO21]. Uniform [Teh16]. Uniqueness [IR21]. Upgrade [CL21b]. Use [RW12]. Using [AN15, BAGM20, BYY12, CASB22, CM12, DB20, Hep10, PO10, ZRA14]. Utilities [FMM11a, Rás15]. Utility [AD23, BD21, BCDD21, BMZ11, BCX19, BS13b, CL21a, DM22, DZ23, EM21, GLZ20, HKZ17, KQY22, KS10, LLX19, MWZZ22, MW13, Mos21, NM13, VF16, WX21, Zho21, dRP22]. Utilizing [PBH10].

Valuation [AC17, BKX12, BR15, BPO18, CKN18, FO11, FT20, GS21, GHK21, JS11, Kar15, LW15, Rog10, War13]. Value [AHO20, BFS11, BMZ11, CCFK20, CSS20, For20, LSWY18, MT17, PVW17, YZZ23]. Value-at-Risk [LSWY18]. Valued [HH10, Hep10, RX17]. Vanillas [DHL15]. Vanna [ARS21]. Varadhan [DF18]. Variables [BKMMS21]. Variance [ATL12, ALMY23, BFS11, BHY19, CLL21, CCW19, GV16, GKR14, HSX15, HX16, HSX23, LLLY18, LLCMA16, MPR14, NPS17, Sch15, SZ22a, SZ22b, vSDF19, vSDF21]. Variate [CMR19]. Variation [CL23, GKR14, vSDF19]. Variational [BK18, CCY12, LT19, NJZB23]. Various [BPO18]. Varying [BCMS22, GS20]. Vector [FPR^{+}18]. Versus [AKK13, LN22, EE22]. via [BDN15, BEV14, BMMB18, BCC15, CMN17, CCC^{+}21, CLP18, LZ17, LLM^{+}23, Mon13a, Qi23, Tia23]. Views [CFRT16]. Viscosity [BCM10, IR21]. VIX [AG20, AGLG22, DHL15, GLOW22, Guy22, Pap18]. Vol [Aly14]. Vol-of-Vol [Aly14]. Volatilities [GV15]. Volatility [AE19, AG20, AKT19, Alm12, AL17, ARS21, AGLG22, Aly14, AFLZ17, BLP16, BCM10, BHP21, BQY22, BFN22, BKX12, BFS11, BK18, CMN17, CM10, CC12, CFR13, CCC^{+}21, CCSW19, CMP21, CMR18, CMR19, CKN18, DHJ17, DF18, De 21, DJ23, EJM21, EFGR19, FLGL18, FJL12, FZ17, FL11, FJL11, FR14, FN18, GS17, GS20, Gas23, GG14, GK10, Gul10, Gul18, GJMN16, HK17a, HK19, HW21, HLLR16, HJT20, JR15, JKR18, JT20, JO23, JT11a, JT11b, LS21, Liu15, LS16, MKPS12, NPS17, NW21, NJZB23, ÖH23, PW22, PZ16, Pun21, Sap20, Teh16, FFF10]. Volterra [AD23, CPZ22, FZ23, Gul18, JMP21]. Volume [CJ16]. VWAP [GR14].

Wages [BGZ22]. Walk [LP21].

Wasserstein [BW23]. Watermark [JLS20]. Wavelet [HK18]. Wavelet-Based [HK18]. Way [RW12]. Weak [BFN22, CD12, Gas23, HK17b, HSZ17, KHOL10]. Wealth [vSDF21, KP10a]. Weighted [CJ16, HSX15, LLP23]. Weights [TY12]. Well [NW21]. Well-Posedness [NW21]. while [BPY16]. Wiener [FTT10, Lel18]. Wind [TT18]. Winning [Cot21]. Wise [Zha23]. Wishart [AKT19]. Withdrawal [HZK17]. Withdrawals [HZK17]. Without [MR18, VF16, Yam22]. Worst [LSWY18]. Worst-Case [LSWY18].

xVA [BGO21].

Yield [FT15].

Zero [ARS21, DHJ17].

References

Albrecher:2022:ORD

[AAM22] Hansjörg Albrecher, Pablo Azcue, and Nora Muler. Optimal ratcheting of dividends in a Brownian risk model. SIAM Journal on Financial Mathematics, 13(3):657–701, ???? 2022. CODEN SJFMBJ. ISSN 1945-

497X. URL https://epubs.
siam.org/doi/10.1137/20M1387171.

Acharya:2021:ROP

[ABRR21] Subas Acharya, Alain Bensoussan, Dmitrii Rachinskii, and Alejandro Rivera. Real options problem with nonsmooth obstacle. SIAM Journal on Financial Mathematics, 12(4):1508–1552, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Angoshtari:2019:ODD

[ABY19] Bahman Angoshtari, Erhan Bayraktar, and Virginia R. Young. Optimal dividend distribution under drawdown and ratcheting constraints on dividend rates. SIAM Journal on Financial Mathematics, 10(2): 547–577, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Angoshtari:2022:OIC

[ABY22] Bahman Angoshtari, Erhan Bayraktar, and Virginia R. Young. Optimal investment and consumption under a habit-formation constraint. SIAM Journal on Financial Mathematics, 13(1):321– 352, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1397891.

Angoshtari:2023:OCU

[ABY23] Bahman Angoshtari, Erhan Bayraktar, and Virginia R. Young. Optimal consumption under a habit-formation constraint: The deterministic case. SIAM Journal on Financial Mathematics, 14(2):557– 597, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/22M1471560.

Armenti:2017:CCV

[AC17] Yannick Armenti and Stéphane Crépey. Central clearing valuation adjustment. SIAM Journal on Financial Mathematics, 8(1): 274–313, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Armenti:2018:MSR

[ACDP18] Yannick Armenti, Stéphane Crépey, Samuel Drapeau, and Antonis Papapantoleon. Multivariate shortfall risk allocation and systemic risk. SIAM Journal on Financial Mathematics, 9 (1):90–126, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Altay:2019:POL

[ACE19] Sühan Altay, Katia Colaneri, and Zehra Eksi. Portfolio optimization for a large investor controlling market sentiment under partial information. SIAM Journal on Financial Mathematics, 10(2):512–546, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Angeris:2023:SCP

[ACEL23] Guillermo Angeris, Tarun Chitra, Alex Evans, and Matthew Lorig. Short communication: A primer on perpetuals. SIAM Journal on Financial Mathematics, 14(1):SC17–SC30, March 2023. CODEN SJFMBJ. ISSN 1945-497X.

Aid:2014:PNM

[ACLP14] René Aïd, Luciano Campi, Nicolas Langrené, and Huyên Pham. A probabilistic numerical method for optimal multiple switching problems in high dimension. SIAM Journal on Financial Mathematics, 5(1): 191–231, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Achtsis:2013:CSB

[ACN13] Nico Achtsis, Ronald Cools, and Dirk Nuyens. Conditional sampling for barrier option pricing under the LT method. SIAM Journal on Financial Mathematics, 4(1):327–352, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Aichinger:2023:UMM

[AD23] Florian Aichinger and Sascha Desmettre. Utility maximization in multivariate Volterra models. SIAM Journal on Financial Mathematics, 14(1): 52–98, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1464543.

AbiJaber:2019:MAR

[AE19] Eduardo Abi Jaber and Omar El Euch. Multifactor approximation of rough volatility models. SIAM Journal on Financial Mathematics, 10(2):309– 349, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Azimzadeh:2015:EOB

[AF15] P. Azimzadeh and P. A. Forsyth. The existence of optimal bangbang controls for GMxB contracts. SIAM Journal on Financial Mathematics, 6(1):117–139, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Armstrong:2017:STA

[AFLZ17] John Armstrong, Martin Forde, Matthew Lorig, and Hongzhong Zhang. Small-time asymptotics under local-stochastic volatility with a jump-to-default: Curvature and the heat kernel expansion. SIAM Journal on Financial Mathematics, 8(1):82–113, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Amini:2020:SRN

[AFM20] Hamed Amini, Damir Filipović, and Andreea Minca. Systemic risk in networks with a central node. SIAM Journal on Financial Mathematics, 11(1):60–98, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Acciaio:2020:SCI

[AG20] Beatrice Acciaio and Julien Guyon. Short communication: Inversion of convex ordering:

Local volatility does not maximize the price of VIX futures. SIAM Journal on Financial Mathematics, 11(1):SC1–SC13, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Alos:2022:SPV

[AGLG22] Elisa Alòs, David García-Lorite, and Aitor Muguruza Gonzalez. On smile properties of volatility derivatives: Understanding the VIX skew. SIAM Journal on Financial Mathematics, 13 (1):32–69, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/19M1269981.

Anthropelos:2022:CFM

[AGZ22] Michail Anthropelos, Tianran Geng, and Thaleia Zariphopoulou. Competition in fund management and forward relative performance criteria. SIAM Journal on Financial Mathematics, 13(4):1271– 1301, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/20M1376169.

Ahn:2015:CPO

[AHJ15] Andrew Ahn, Martin Haugh, and Ashish Jain. Consistent pricing of options on leveraged ETFs. SIAM Journal on Financial Mathematics, 6(1):559–593, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Aksamit:2020:RFQ

[AHO20] Anna Aksamit, Zhaoxu Hou, and Jan Obłój. Robust framework for quantifying the value of information in pricing and hedging. SIAM Journal on Financial Mathematics, 11(1):27–59, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Avellaneda:2022:PEU

[AHPP22] Marco Avellaneda, Brian Healy, Andrew Papanicolaou, and George Papanicolaou. Principal eigenportfolios for U.S. equities. SIAM Journal on Financial Mathematics, 13(3):702– 744, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/20M1383501.

Abad:2015:PSM

[AI15] Carlos Abad and Garud Iyengar. Portfolio selection with multiple spectral risk constraints. SIAM Journal on Financial Mathematics, 6(1):467–486, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Abergel:2015:LTB

[AJ15] Frédéric Abergel and Aymen Jedidi. Long-time behavior of a Hawkes process-based limit order book. SIAM Journal on Financial Mathematics, 6(1):1026– 1043, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Ankirchner:2013:HFP

[AKK13] Stefan Ankirchner, Peter Kratz, and Thomas Kruse. Hedging forward positions: Basis risk versus liquidity costs. SIAM Journal on Financial Mathematics, 4(1):

668–696, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Alfonsi:2019:LTL

[AKT19] Aurélien Alfonsi, David Krief, and Peter Tankov. Long-time large deviations for the multiasset Wishart stochastic volatility model and option pricing. SIAM Journal on Financial Mathematics, 10(4):942–976, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Ackermann:2021:OTE

[AKU21] Julia Ackermann, Thomas Kruse, and Mikhail Urusov. Optimal trade execution in an order book model with stochastic liquidity parameters. SIAM Journal on Financial Mathematics, 12(2):788–822, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Alos:2017:CSS

[AL17] Elisa Alòs and Jorge A. León.

On the curvature of the smile in stochastic volatility models. SIAM Journal on Financial Mathematics, 8(1):373–399, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Almgren:2012:OTS

[Alm12] Robert Almgren. Optimal trading with stochastic liquidity and volatility. SIAM Journal on Financial Mathematics, 3(1): 163–181, ???? 2012. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v3/i1/p163_s1.

Azcue:2023:ORM

[ALMY23] Pablo Azcue, Xiaoqing Liang, Nora Muler, and Virginia R. Young. Optimal reinsurance to minimize the probability of drawdown under the mean-variance premium principle: Asymptotic analysis. SIAM Journal on Financial Mathematics, 14(1):279– 313, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1461666.

Aly:2014:OPS

[Aly14] S. M. Ould Aly. Option pricing for stochastic volatility models: Vol-of-vol expansion. SIAM Journal on Financial Mathematics, 5(1):729–752, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Abergel:2011:NLR

[AM11] Frédéric Abergel and Nicolas Millot. Nonquadratic local riskminimization for hedging contingent claims in incomplete markets. SIAM Journal on Financial Mathematics, 2(1): 342–356, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL http://epubs.
                                        siam.org/sifin/resource/1/
                                        sjfmbj/v2/i1/p342_s1.

Amini:2015:CIC

[AMS15] Hamed Amini, Andreea Minca, and Agnès Sulem. Control of interbank contagion under partial information. SIAM Journal on Financial Mathematics, 6(1):

1195–1219, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Altmayer:2015:MMC

[AN15] Martin Altmayer and Andreas Neuenkirch. Multilevel Monte Carlo quadrature of discontinuous payoffs in the generalized Heston model using Malliavin integration by parts. SIAM Journal on Financial Mathematics, 6(1):22–52, ???? 2015. CO-DEN SJFMBJ. ISSN 1945-497X.

Anthropelos:2014:FEP

[Ant14] Michail Anthropelos. Forward exponential performances: Pricing and optimal risk sharing. SIAM Journal on Financial Mathematics, 5(1):626–655, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Alvarez:2023:OBC

[ANW23] Guillermo Alonso Alvarez, Sergey Nadtochiy, and Kevin Webster. Optimal brokerage contracts in almgren-chriss model with multiple clients. SIAM Journal on Financial Mathematics, 14(3): 855–878, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/22M1490156.

Arai:2011:GDB

[Ara11] Takuji Arai. Good deal bounds induced by shortfall risk. SIAM Journal on Financial Mathematics, 2(1):1–21, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.

                                        siam.org/sifin/resource/1/
                                        sjfmbj/v2/i1/p1_s1.

Arai:2014:CRM

[Ara14] Takuji Arai. Convex risk measures for Càdlàg processes on Orlicz hearts. SIAM Journal on Financial Mathematics, 5(1): 609–625, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Armstrong:2018:MC

[Arm18] John Armstrong. The Markowitz category. SIAM Journal on Financial Mathematics, 9(3):994– 1016, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Alos:2021:DBV

[ARS21] Elisa Alòs, Frido Rolloos, and Kenichiro Shiraya. On the difference between the volatility swap strike and the zero vanna implied volatility. SIAM Journal on Financial Mathematics, 12 (2):690–723, ???? 2021. CO-DEN SJFMBJ. ISSN 1945-497X.

Alfonsi:2010:OTE

[AS10] Aurélien Alfonsi and Alexander Schied. Optimal trade execution and absence of price manipulations in limit order book models. SIAM Journal on Financial Mathematics, 1(1):490–522, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Agarwal:2018:PBU

[AS18] Ankush Agarwal and Ronnie Sircar. Portfolio benchmarking under drawdown constraint and stochastic Sharpe ratio. SIAM Journal on Financial Mathematics, 9(2):435–464, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Avanesyan:2022:PMF

[AS22] Levon Avanesyan and Ronnie Sircar. Power mixture forward performance processes. SIAM Journal on Financial Mathematics, 13(3):1040– 1062, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/20M1385500.

Alfonsi:2012:OBR

[ASS12] Aurélien Alfonsi, Alexander Schied, and Alla Slynko. Order book resilience, price manipulation, and the positive portfolio problem. SIAM Journal on Financial Mathematics, 3(1): 511–533, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Abbas-Turki:2012:AOM

[ATL12] L. A. Abbas-Turki and B. Lapeyre.

American options by Malliavin calculus and nonparametric variance and bias reduction methods. SIAM Journal on Financial Mathematics, 3(1):479–510, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Avellaneda:2010:PDL

[AZ10] Marco Avellaneda and Stanley Zhang. Path-dependence of leveraged ETF returns. SIAM Journal on Financial Mathematics, 1(1):586–603, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Begin:2020:EJD

[BAGM20] Jean-François Bégin, Diego Amaya, Geneviève Gauthier, and Marie-Ève Malette. On the estimation of jump-diffusion models using intraday data: a filtering-based approach. SIAM Journal on Financial Mathematics, 11(4):1168–1208, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Bouchard:2016:BDR

[BBC16] Bruno Bouchard, Géraldine Bouveret, and Jean-François Chassagneux. A backward dual representation for the quantile hedging of Bermudan options. SIAM Journal on Financial Mathematics, 7(1):215–235, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Bartesaghi:2020:RDC

[BBC^{+}20] Paolo Bartesaghi, Michele Benzi, Gian Paolo Clemente, Rosanna Grassi, and Ernesto Estrada. Risk-dependent centrality in economic and financial networks. SIAM Journal on Financial Mathematics, 11(2):526– 565, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Bayer:2021:ROS

[BBH^{+}21] Christian Bayer, Denis Belomestny, Paul Hager, Paolo Pigato, and John Schoenmakers. Randomized optimal stopping algorithms and their convergence analysis. SIAM Journal on Financial Mathematics, 12(3):1201–1225, ???? 2021.

CODEN SJFMBJ. ISSN 1945-497X.

Bauer:2012:MFS

[BBK12] Daniel Bauer, Fred Espen Benth, and Rüdiger Kiesel. Modeling the forward surface of mortality. SIAM Journal on Financial Mathematics, 3(1): 639–666, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Bo:2015:SRI

[BC15] Lijun Bo and Agostino Capponi. Systemic risk in interbanking networks. SIAM Journal on Financial Mathematics, 6(1): 386–424, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Blanchard:2022:SCS

[BC22] Romain Blanchard and Laurence Carassus. Short communication: Super-replication prices with multiple priors in discrete time. SIAM Journal on Financial Mathematics, 13(2):SC53– SC65, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/22M1470013.

Bielecki:2015:DCF

[BCC15] Tomasz R. Bielecki, Igor Cialenco, and Tao Chen. Dynamic conic finance via backward stochastic difference equations. SIAM Journal on Financial Mathematics, 6(1):1068– 1122, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Bayraktar:2021:SCN

[BCDD21] Erhan Bayraktar, Christoph Czichowsky, Leonid Dolinskyi, and Yan Dolinsky. Short communication: a note on utility maximization with proportional transaction costs and stability of optimal portfolios. SIAM Journal on Financial Mathematics, 12(4):SC115–SC125, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Bardi:2010:CVM

[BCM10] Martino Bardi, Annalisa Cesaroni, and Luigi Manca. Convergence by viscosity methods in multiscale financial models with stochastic volatility. SIAM Journal on Financial Mathematics, 1(1):230–265, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Belak:2022:OIT

[BCMS22] Christoph Belak, An Chen, Carla Mereu, and Robert Stelzer. Optimal investment with time-varying stochastic endowments. SIAM Journal on Financial Mathematics, 13(3): 969–1003, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1453402.

Bayraktar:2023:NNA

[BCN23] Erhan Bayraktar, Asaf Cohen, and April Nellis. A neural network approach to highdimensional optimal switching problems with jumps in energy markets. SIAM Journal on Financial Mathematics, 14(4): 1028–1061, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/22M1527246.

Benezet:2021:NSQ

[BCR21] Cyril Bénézet, Jean-François Chassagneux, and Christoph Reisinger. A numerical scheme for the quantile hedging problem. SIAM Journal on Financial Mathematics, 12(1):110– 157, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Basei:2014:OES

[BCV14] Matteo Basei, Annalisa Cesaroni, and Tiziano Vargiolu. Optimal exercise of swing contracts in energy markets: an integral constrained stochastic optimal control problem. SIAM Journal on Financial Mathematics, 5(1): 581–608, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Bian:2019:UMU

[BCX19] Baojun Bian, Xinfu Chen, and Zuo Quan Xu. Utility maximization under trading constraints with discontinuous utility. SIAM Journal on Financial Mathematics, 10(1):243–260, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Beissner:2018:DGE

[BD18] Patrick Beissner and Laurent Denis. Duality and general equilibrium theory under Knightian uncertainty. SIAM Journal on Financial Mathematics, 9(1):

381–400, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Bauerle:2020:POF

[BD20] Nicole Bäuerle and Sascha Desmettre. Portfolio optimization in fractional and rough Heston models. SIAM Journal on Financial Mathematics, 11(1): 240–273, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Bank:2021:SCN

[BD21] Peter Bank and Yan Dolinsky. Short communication: a note on utility indifference pricing with delayed information. SIAM Journal on Financial Mathematics, 12(2):SC31–SC43, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Bayraktar:2018:RTA

[BDG18] Erhan Bayraktar, Yan Dolinsky, and Jia Guo. Recombining tree approximations for optimal stopping for diffusions. SIAM Journal on Financial Mathematics, 9(2):602–633, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Bergault:2022:MAO

[BDG22] Philippe Bergault, Fayçal Drissi, and Olivier Guéant. Multiasset optimal execution and statistical arbitrage strategies under Ornstein–Uhlenbeck dynamics. SIAM Journal on Financial Mathematics, 13(1):353– 390, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1407756.

Bensoussan:2010:ROG

[BDH10] Alain Bensoussan, J. David Diltz, and SingRu Hoe. Real options games in complete and incomplete markets with several decision makers. SIAM Journal on Financial Mathematics, 1(1): 666–728, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Bouchard:2011:OCT

[BDL11] Bruno Bouchard, Ngoc-Minh Dang, and Charles-Albert Lehalle. Optimal control of trading algorithms: a general impulse control approach. SIAM Journal on Financial Mathematics, 2 (1):404–438, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p404_s1.

Belomestny:2015:PBO

[BDN15] Denis Belomestny, Fabian Dickmann, and Tigran Nagapetyan. Pricing Bermudan options via multilevel approximation methods. SIAM Journal on Financial Mathematics, 6(1):448–466, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Bender:2011:PDP

[Ben11]  Christian Bender.  Primal

and dual pricing of multiple exercise options in continuous time. SIAM Journal on Financial Mathematics, 2(1): 562–586, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.

siam.org/sifin/resource/1/
sjfmbj/v2/i1/p562_s1.

Bayer:2023:PHD

[BEST23] Christian Bayer, Martin Eigel, Leon Sallandt, and Philipp Trunschke. Pricing highdimensional Bermudan options with hierarchical tensor formats. SIAM Journal on Financial Mathematics, 14(2):383– 406, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1402170.

Benth:2014:ALS

[BEV14] Fred Espen Benth, Heidar Eyjolfsson, and Almut E. D. Veraart. Approximating Lévy semistationary processes via Fourier methods in the context of power markets. SIAM Journal on Financial Mathematics, 5 (1):71–98, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Bank:2014:OOS

[BF14a] Peter Bank and Antje Fruth. Optimal order scheduling for deterministic liquidity patterns. SIAM Journal on Financial Mathematics, 5(1):137–152, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Bressan:2014:DBS

[BF14b] Alberto Bressan and Giancarlo Facchi. Discrete bidding strategies for a random incoming order. SIAM Journal on Financial Mathematics, 5(1):50–70, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Bichuch:2019:OFS

[BF19] Maxim Bichuch and Zachary Feinstein. Optimization of fire sales and borrowing in systemic risk. SIAM Journal on Financial Mathematics, 10(1):68–88, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Bayer:2022:SCW

[BFN22] Christian Bayer, Masaaki Fukasawa, and Shonosuke Nakahara. Short communication: On the weak convergence rate in the discretization of rough volatility models. SIAM Journal on Financial Mathematics, 13(2):SC66– SC73, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/22M1482871.

Beiglbock:2011:MVO

[BFS11] Mathias Beiglböck, Peter Friz, and Stephan Sturm. Is the minimum value of an option on variance generated by local volatility? SIAM Journal on Financial Mathematics, 2(1):213–220, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p213_s1.

Burzoni:2021:SCR

[BFZ21] Matteo Burzoni, Marco Frittelli, and Federico Zorzi. Short communication: Robust marketadjusted systemic risk measures. SIAM Journal on Financial Mathematics, 12(3):SC70–SC82, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Bolker:2021:SCS

[BGH21] Benjamin M. Bolker, Matheus R. Grasselli, and Emma Holmes. Short communication: Sensitivity analysis of an integrated climate-economic model. SIAM Journal on Financial Mathematics, 12(2):SC44–SC57, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Benhamou:2010:TDH

[BGM10] E. Benhamou, E. Gobet, and M. Miri. Time dependent Heston model. SIAM Journal on Financial Mathematics, 1(1):289– 325, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Biagini:2021:UAX

[BGO21] Francesca Biagini, Alessandro Gnoatto, and Immacolata Oliva. A unified approach to xVA with CSA discounting and initial margin. SIAM Journal on Financial Mathematics, 12(3): 1013–1053, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Bourgey:2020:MLC

[BGR20] Florian Bourgey, Emmanuel Gobet, and Clément Rey. Metamodel of a large credit risk portfolio in the Gaussian copula model. SIAM Journal on Financial Mathematics, 11(4): 1098–1136, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Bernis:2023:IRT

[BGSS23] Guillaume Bernis, Matthieu Garcin, Simone Scotti, and Carlo Sgarra. Interest rates term structure models driven by Hawkes processes. SIAM Journal on Financial Mathematics, 14(4):1062–1079, ???? 2023. CODEN SJFMBJ. ISSN 1945-

497X. URL https://epubs.
siam.org/doi/10.1137/22M1502604.

Biagini:2022:RPC

[BGZ22] Sara Biagini, Fausto Gozzi, and Margherita Zanella. Robust portfolio choice with sticky wages. SIAM Journal on Financial Mathematics, 13(3):1004– 1039, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1429722.

Backhoff:2016:CAP

[BH16] Julio Backhoff and Ulrich Horst.

Conditional analysis and a principal-agent problem. SIAM Journal on Financial Mathematics, 7(1):477–507, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Bush:2011:SEE

[BHH^{+}11] N. Bush, B. M. Hambly, H. Haworth, L. Jin, and C. Reisinger. Stochastic evolution equations in portfolio credit modelling. SIAM Journal on Financial Mathematics, 2(1):627–664, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p627_s1.

Brody:2020:TCI

[BHM20] Dorje Brody, Lane Hughston, and Bernhard Meister. Theory of cryptocurrency interest rates. SIAM Journal on Financial Mathematics, 11(1):148– 168, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Bayer:2021:LMR

[BHP21] Christian Bayer, Fabian A. Harang, and Paolo Pigato. Logmodulated rough stochastic volatility models. SIAM Journal on Financial Mathematics, 12(3):1257–1284, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Burkovska:2015:RBM

[BHSW15] O. Burkovska, B. Haasdonk, J. Salomon, and B. Wohlmuth. Reduced basis methods for pricing options with the Black– Scholes and Heston models. SIAM Journal on Financial Mathematics, 6(1):685–712, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Belomestny:2018:RBC

[BHU18] Denis Belomestny, Stefan Häfner, and Mikhail Urusov. Regressionbased complexity reduction of the nested Monte Carlo methods. SIAM Journal on Financial Mathematics, 9(2):665–689, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Bensoussan:2019:MVA

[BHY19] Alain Bensoussan, SingRu Celine Hoe, and Zhongfeng Yan.

A mean-variance approach to capital investment optimization. SIAM Journal on Financial Mathematics, 10(1):156– 180, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Bayraktar:2015:HAO

[BHZ15] Erhan Bayraktar, Yu-Jui Huang, and Zhou Zhou. On hedging American options under model uncertainty. SIAM Journal on Financial Mathematics, 6(1): 425–447, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Bichuch:2012:AAO

[Bic12] Maxim Bichuch. Asymptotic analysis for optimal investment in finite time with transaction costs. SIAM Journal on Financial Mathematics, 3(1):433–458, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Beveridge:2012:ISD

[BJ12] Christopher Beveridge and Mark Joshi. Interpolation schemes in the displaced-diffusion LIBOR market model. SIAM Journal on Financial Mathematics, 3(1): 593–604, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Benth:2015:DPE

[BK15] Fred Espen Benth and Paul Krühner. Derivatives pricing in energy markets: an infinitedimensional approach. SIAM Journal on Financial Mathematics, 6(1):825–869, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Bonnans:2018:VAO

[BK18] J. Frédéric Bonnans and Axel Kröner. Variational analysis for options with stochastic volatility and multiple factors. SIAM Journal on Financial Mathematics, 9(2):465–492, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Bank:2022:MOI

[BK22] Peter Bank and Laura Körber. Merton’s optimal investment problem with jump signals. SIAM Journal on Financial Mathematics, 13(4):1302–1325, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X. URL https:/

                                        /epubs.siam.org/doi/10.1137/
                                        21M1450161.

Bellini:2021:LIF

[BKMMS21] Fabio Bellini, Pablo Koch-Medina, Cosimo Munari, and Gregor Svindland. Lawinvariant functionals on general spaces of random variables. SIAM Journal on Financial Mathematics, 12(1):318– 341, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Bayraktar:2012:VES

[BKX12] Erhan Bayraktar, Constantinos Kardaras, and Hao Xing. Valuation equations for stochastic volatility models. SIAM Journal on Financial Mathematics, 3(1): 351–373, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Bernard:2013:PHC

[BL13] Carole Bernard and Wenbo V. Li. Pricing and hedging of Cliquet options and locally capped contracts. SIAM Journal on Financial Mathematics, 4(1):353– 371, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Bechler:2015:OED

[BL15] Kyle Bechler and Michael Ludkovski. Optimal execution with dynamic order flow imbalance. SIAM Journal on Financial Mathematics, 6(1):1123– 1151, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Bouselmi:2016:CPA

[BL16] Aych Bouselmi and Damien Lamberton. The critical price of the American put near maturity in the jump diffusion model. SIAM Journal on Financial Mathematics, 7(1):236–272, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Benth:2021:CPP

[BL21] Fred Espen Benth and Silvia Lavagnini. Correlators of polynomial processes. SIAM Journal on Financial Mathematics, 12(4):1374–1415, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Bielagk:2017:EPU

[BLD17] Jana Bielagk, Arnaud Lionnet, and Gonçalo Dos Reis. Equilibrium pricing under relative performance concerns. SIAM Journal on Financial Mathematics, 8(1):435–482, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Baltean-Lugojan:2016:RNC

[BLP16] Radu Baltean-Lugojan and Panos Parpas. Robust numerical calibration for implied volatility expansion models. SIAM Journal on Financial Mathematics, 7(1):917–946, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Biagini:2018:LIA

[BMMB18] Francesca Biagini, Andrea Mazzon, and Thilo Meyer-Brandis. Liquidity induced asset bubbles via flows of ELMMs. SIAM Journal on Financial Mathematics, 9(2):800–834, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Biagini:2019:FAB

[BMMB19] Francesca Biagini, Andrea Mazzon, and Thilo Meyer-Brandis. Financial asset bubbles in banking networks. SIAM Journal on Financial Mathematics, 10(2): 430–465, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Bressan:2017:SMO

[BMNP17] Alberto Bressan, Antonio Marigonda, Khai T. Nguyen, and Michele Palladino. A stochastic model of optimal debt management and bankruptcy. SIAM Journal on Financial Mathematics, 8(1): 841–873, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Bouchard:2016:HUE

[BMS16] Bruno Bouchard, Ludovic Moreau, and H. Mete Soner. Hedging under an expected loss constraint with small transaction costs. SIAM Journal on Financial Mathematics, 7(1):508–551, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Bian:2011:SVF

[BMZ11] Baojun Bian, Sheng Miao, and Harry Zheng. Smooth value functions for a class of nonsmooth utility maximization problems. SIAM Journal on Financial Mathematics, 2(1): 727–747, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p727_s1.

Biagini:2015:FFB

[BN15] Francesca Biagini and Sorin Nedelcu. The formation of financial bubbles in defaultable markets. SIAM Journal on Financial Mathematics, 6(1):530–558, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Bion-Nadal:2020:FDR

[BND20] Jocelyne Bion-Nadal and Giulia Di Nunno. Fully-dynamic risk-indifference pricing and nogood-deal bounds. SIAM Journal on Financial Mathematics, 11(2):620–658, ???? 2020. CO-DEN SJFMBJ. ISSN 1945-497X.

Benth:2014:PME

[BOL14] Fred Espen Benth and Salvador Ortiz-Latorre. A pricing measure to explain the risk premium in power markets. SIAM Journal on Financial Mathematics, 5(1):

685–728, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Biagini:2013:BGL

[BP13] Sara Biagini and Mustafa Ç. Pinar. The best gain-loss ratio is a poor performance measure. SIAM Journal on Financial Mathematics, 4(1):228–242, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Bellini:2022:SCA

[BP22] Fabio Bellini and Ilaria Peri. Short communication: An axiomatization of Λ-quantiles. SIAM Journal on Financial Mathematics, 13(1):SC26–SC38, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X. URL https:/

                                        /epubs.siam.org/doi/10.1137/
                                        21M1444278.

Borovykh:2018:ECV

[BPO18] Anastasia Borovykh, Andrea Pascucci, and Cornelis W. Oosterlee. Efficient computation of various valuation adjustments under local Lévy models. SIAM Journal on Financial Mathematics, 9(1):251–273, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Baldacci:2021:OMT

[BPR21] Bastien Baldacci, Dylan Possamaï, and Mathieu Rosenbaum. Optimal make-take fees in a multi market-maker environment. SIAM Journal on Financial Mathematics, 12(1):446– 486, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Bayraktar:2016:PTL

[BPY16] Erhan Bayraktar, S. David Promislow, and Virginia R. Young. Purchasing term life insurance to reach a bequest goal while consuming. SIAM Journal on Financial Mathematics, 7(1): 183–214, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Bayer:2022:POU

[BQY22] Christian Bayer, Jinniao Qiu, and Yao Yao. Pricing options under rough volatility with backward SPDEs. SIAM Journal on Financial Mathematics, 13(1): 179–212, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/20M1357639.

Bielecki:2015:VHC

[BR15] Tomasz R. Bielecki and Marek Rutkowski. Valuation and hedging of contracts with funding costs and collateralization. SIAM Journal on Financial Mathematics, 6(1):594–655, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Biagini:2013:RML

[BS13a] Francesca Biagini and Irene Schreiber. Risk-minimization for life insurance liabilities. SIAM Journal on Financial Mathematics, 4(1):243–264, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Bichuch:2013:UMT

[BS13b] Maxim Bichuch and Steven Shreve. Utility maximization trading two futures with transaction costs. SIAM Journal on Financial Mathematics, 4(1): 26–85, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Bosserhoff:2023:RDH

[BS23] Frank Bosserhoff and Mitja Stadje. Robustness of delta hedging in a jump-diffusion model. SIAM Journal on Financial Mathematics, 14(2):663– 703, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/22M149435X.

BenTahar:2010:MPT

[BST10] Imen Ben Tahar, H. Mete Soner, and Nizar Touzi. Merton problem with taxes: Characterization, computation, and approximation. SIAM Journal on Financial Mathematics, 1(1):366– 395, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Bernhart:2011:FDA

[BTW11] Marie Bernhart, Peter Tankov, and Xavier Warin. A finitedimensional approximation for pricing moving average options. SIAM Journal on Financial Mathematics, 2(1):989– 1013, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p989_s1.

Burzoni:2016:AHM

[Bur16] Matteo Burzoni. Arbitrage and hedging in model-independent markets with frictions. SIAM Journal on Financial Mathematics, 7(1):812–844, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Bauerle:2012:RIP

[BUV12] Nicole Bäuerle, Sebastian P. Urban, and Luitgard A. M. Veraart. The relaxed investor with partial information. SIAM Journal on Financial Mathematics, 3 (1):304–327, ???? 2012. CO-DEN SJFMBJ. ISSN 1945-497X.

Bank:2019:OIT

[BV19] Peter Bank and Moritz Voß. Optimal investment with transient price impact. SIAM Journal on Financial Mathematics, 10(3): 723–768, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Broden:2011:CHO

[BW11] Mats Brodén and Magnus Wiktorsson. On the convergence of higher order hedging schemes: The delta–gamma case. SIAM Journal on Financial Mathematics, 2(1):55–78, ???? 2011. CODEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p55_s1.

Bartl:2023:SMO

[BW23] Daniel Bartl and Johannes Wiesel. Sensitivity of multiperiod optimization problems with respect to the adapted Wasserstein distance. SIAM Journal on Financial Mathematics, 14(2):704–720, ???? 2023.

CODEN SJFMBJ. ISSN 1945-

                                        497X. URL https://epubs.
                                        siam.org/doi/10.1137/22M1537746.

Bensoussan:2014:TCP

[BWYY14] A. Bensoussan, K. C. Wong, S. C. P. Yam, and S. P. Yung. Time-consistent portfolio selection under short-selling prohibition: From discrete to continuous setting. SIAM Journal on Financial Mathematics, 5(1): 153–190, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Bayraktar:2022:SCS

[BWZ22] Erhan Bayraktar, Zhenhua Wang, and Zhou Zhou. Short communication: Stability of time-inconsistent stopping for one-dimensional diffusions. SIAM Journal on Financial Mathematics, 13(4):SC123–SC135, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X. URL https://

                                        epubs.siam.org/doi/10.1137/
                                        22M1510005.

Bensoussan:2012:TTP

[BYY12] Alain Bensoussan, ZhongFeng Yan, and G. Yin. Thresholdtype policies for real options using regime-switching models. SIAM Journal on Financial Mathematics, 3(1):667–689, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Bayraktar:2014:CSP

[BZ14] Erhan Bayraktar and Zhou Zhou. On controller-stopper problems with jumps and their applications to indifference pricing of American options. SIAM Journal on Financial Mathematics, 5(1):20–49, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Barski:2020:CEG

[BZ20] Michał Barski and Jerzy Zabczyk.

On CIR equations with general factors. SIAM Journal on Financial Mathematics, 11(1): 131–147, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Bayraktar:2019:TCS

[BZZ19] Erhan Bayraktar, Jingjie Zhang, and Zhou Zhou. Time consistent stopping for the meanstandard deviation problem — the discrete time case. SIAM Journal on Financial Mathematics, 10(3):667–697, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Cartea:2022:DES

[CASB22] Álvaro Cartea, Imanol Pérez Arribas, and Leandro Sánchez-Betancourt. Double-execution strategies using path signatures. SIAM Journal on Financial Mathematics, 13(4):1379– 1417, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1456467.

Carr:2012:ECM

[CC12] Peter Carr and Laurent Cousot.

Explicit constructions of martingales calibrated to given implied volatility smiles. SIAM Journal on Financial Mathematics, 3(1):182–214, ???? 2012.

CODEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v3/i1/p182_s1.

Caravenna:2016:GSA

[CC16] Francesco Caravenna and Jacopo Corbetta. General smile asymptotics with bounded maturity. SIAM Journal on Financial Mathematics, 7(1):720–759, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Chataigner:2021:SCB

[CCC^{+}21] Marc Chataigner, Areski Cousin, Stéphane Crépey, Matthew Dixon, and Djibril Gueye. Short communication: Beyond surrogate modeling: Learning the local volatility via shape constraints. SIAM Journal on Financial Mathematics, 12(3): SC58–SC69, ???? 2021. CO-DEN SJFMBJ. ISSN 1945-497X.

Ceci:2020:VAD

[CCFK20] Claudia Ceci, Katia Colaneri, Rüdiger Frey, and Verena Köck. Value adjustments and dynamic hedging of reinsurance counterparty risk. SIAM Journal on Financial Mathematics, 11(3): 788–814, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Cheng:2011:CFA

[CCL^{+}11] Wen Cheng, Nick Costanzino, John Liechty, Anna Mazzucato, and Victor Nistor. Closed-form asymptotics and numerical approximations of 1D parabolic equations with applications to option pricing. SIAM Journal on Financial Mathematics, 2(1):901–934, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p901_s1.

Chen:2019:SVA

[CCSW19] Kexin Chen, Mei Choi Chiu, Yong Hyun Shin, and Hoi Ying Wong. Stochastic volatility asymptotics for optimal subsistence consumption and investment with bankruptcy. SIAM Journal on Financial Mathematics, 10(4):977–1005, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Chen:2019:TCM

[CCW19] Kexin Chen, Mei Choi Chiu, and Hoi Ying Wong. Timeconsistent mean-variance pairstrading under regime-switching cointegration. SIAM Journal on Financial Mathematics, 10(2): 632–665, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Capponi:2012:VAC

[CCY12] Agostino Capponi, Jaksa Cvitanić, and Türkay Yolcu. A variational approach to contracting under imperfect observations. SIAM Journal on Financial Mathematics, 3(1):605–638, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Chakraborty:2023:ODU

[CCY23] Prakash Chakraborty, Asaf Cohen, and Virginia R. Young. Optimal dividends under model uncertainty. SIAM Journal on Financial Mathematics, 14(2): 497–524, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1447453.

Campi:2012:WIT

[CD12] L. Campi and M. Del Vigna. Weak insider trading and behavioral finance. SIAM Journal on Financial Mathematics, 3(1): 242–279, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Chen:2013:COS

[CD13] Xinfu Chen and Min Dai. Characterization of optimal strategy for multiasset investment and consumption with transaction costs. SIAM Journal on Financial Mathematics, 4(1):857–883, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Coculescu:2022:ITF

[CD22] Delia Coculescu and Aditi Dandapani. Insiders and their free lunches: The role of short positions. SIAM Journal on Financial Mathematics, 13(3):877– 902, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/20M1375826.

Cartea:2017:ATM

[CDJ17] Álvaro Cartea, Ryan Donnelly, and Sebastian Jaimungal. Algorithmic trading with model uncertainty. SIAM Journal on Financial Mathematics, 8(1):635– 671, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Cont:2010:DII

[CDK10] Rama Cont, Romain Deguest, and Yu Hang Kan. Default intensities implied by CDO spreads: Inversion formula and model calibration. SIAM Journal on Financial Mathematics, 1 (1):555–585, ???? 2010. CO-DEN SJFMBJ. ISSN 1945-497X.

Cont:2013:PDM

[CdL13] Rama Cont and Adrien de Larrard. Price dynamics in a Markovian limit order market. SIAM Journal on Financial Mathematics, 4(1):1–25, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Cai:2021:OHP

[CDP21] Cheng Cai, Tiziano De Angelis, and Jan Palczewski. Optimal hedging of a perpetual American put with a single trade. SIAM Journal on Financial Mathematics, 12(2):823–866, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Capponi:2015:DCA

[CF15] Agostino Capponi and Christoph Frei. Dynamic contracting: Accidents lead to nonlinear contracts. SIAM Journal on Financial Mathematics, 6(1):959–983, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Corielli:2010:PAD

[CFP10] Francesco Corielli, Paolo Foschi, and Andrea Pascucci. Parametrix approximation of diffusion transition densities. SIAM Journal on Financial Mathematics, 1(1):833–867, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Carr:2013:WQN

[CFR13] Peter Carr, Travis Fisher, and Johannes Ruf. Why are quadratic normal volatility models analytically tractable? SIAM Journal on Financial Mathematics, 4(1):185–202, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Cai:2016:ODH

[CFRT16] Jiatu Cai, Masaaki Fukasawa, Mathieu Rosenbaum, and Peter Tankov. Optimal discretization of hedging strategies with directional views. SIAM Journal on Financial Mathematics, 7 (1):34–69, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Cartea:2022:OCB

[CFVS22] Álvaro Cartea, Maria Flora, Tiziano Vargiolu, and Georgi Slavov. Optimal cross-border electricity trading. SIAM Journal on Financial Mathematics, 13(1):262–294, ???? 2022. CO-DEN SJFMBJ. ISSN 1945-

497X. URL https://epubs.
siam.org/doi/10.1137/21M1398537.

Calvia:2020:RMP

[CG20] Alessandro Calvia and Emanuela Rosazza Gianin. Risk measures and progressive enlargement of filtration: a BSDE approach. SIAM Journal on Financial Mathematics, 11(3):815–848, ???? 2020.

CODEN SJFMBJ. ISSN 1945-497X.

Cartea:2019:HSR

[CGJ19] Álvaro Cartea, Luhui Gan, and Sebastian Jaimungal. Hedge and speculate: Replicating option payoffs with limit and market orders. SIAM Journal on Financial Mathematics, 10(3): 790–814, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Cuchiero:2023:SBM

[CGSF23] Christa Cuchiero, Guido Gazzani, and Sara Svaluto-Ferro. Signature-based models: Theory and calibration. SIAM Journal on Financial Mathematics, 14 (3):910–957, ???? 2023. CO-DEN SJFMBJ. ISSN 1945-

497X. URL https://epubs.
siam.org/doi/10.1137/22M1512338.

Cartea:2016:CFE

[CJ16] Álvaro Cartea and Sebastian Jaimungal. A closed-form execution strategy to target volume weighted average price. SIAM Journal on Financial Mathematics, 7(1):760–785, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Cartea:2020:TFE

[CJJ20] Álvaro Cartea, Sebastian Jaimungal, and Tianyi Jia. Trading foreign exchange triplets. SIAM Journal on Financial Mathematics, 11(3):690–719, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Chassagneux:2016:EES

[CJM16] Jean-François Chassagneux, Antoine Jacquier, and Ivo Mihaylov. An explicit Euler scheme with strong rate of convergence for financial SDEs with non-Lipschitz coefficients. SIAM Journal on Financial Mathematics, 7(1):993–1021, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Cartea:2016:MUC

[CJQ16] Álvaro Cartea, Sebastian Jaimungal, and Zhen Qin. Model uncertainty in commodity markets. SIAM Journal on Financial Mathematics, 7(1):1–33, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Cartea:2014:BLS

[CJR14] Álvaro Cartea, Sebastian Jaimungal, and Jason Ricci. Buy low, sell high: a high frequency trading perspective. SIAM Journal on Financial Mathematics, 5(1): 415–444, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Cont:2011:DHP

[CK11] Rama Cont and Yu Hang Kan. Dynamic hedging of portfolio credit derivatives. SIAM Journal on Financial Mathematics, 2(1):112–140, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL http://epubs.
                                        siam.org/sifin/resource/1/
                                        sjfmbj/v2/i1/p112_s1.

Chong:2018:CFS

[CK18] Carsten Chong and Claudia Klüppelberg. Contagion in financial systems: a Bayesian network approach. SIAM Journal on Financial Mathematics, 9(1): 28–53, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Cui:2018:GVF

[CKN18] Zhenyu Cui, J. Lars Kirkby, and Duy Nguyen. A general valuation framework for SABR and stochastic local volatility models. SIAM Journal on Financial Mathematics, 9(2):520–563, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Cheridito:2017:DFR

[CKT17] Patrick Cheridito, Michael Kupper, and Ludovic Tangpi. Duality formulas for robust pricing and hedging in discrete time. SIAM Journal on Financial Mathematics, 8(1):738–765, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Chen:2021:MLA

[CL21a] Tao Chen and Michael Ludkovski. A machine learning approach to adaptive robust utility maximization and hedging. SIAM Journal on Financial Mathematics, 12(3):1226–1256, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Chen:2021:FBP

[CL21b] Xinfu Chen and Jin Liang.

A free boundary problem for corporate bond pricing and credit rating under different upgrade and downgrade thresholds. SIAM Journal on Financial Mathematics, 12(3):941– 966, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Carmona:2023:OEQ

[CL23] Rene Carmona and Laura Leal. Optimal execution with quadratic variation inventories. SIAM Journal on Financial Mathematics, 14(3):751– 776, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1416564.

Carr:2021:PVS

[CLL21] Peter Carr, Roger Lee, and Matthew Lorig. Pricing variance swaps on time-changed Markov processes. SIAM Journal on Financial Mathematics, 12(2): 672–689, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Cont:2011:RBO

[CLP11] Rama Cont, Nicolas Lantos, and Olivier Pironneau. A reduced basis for option pricing. SIAM Journal on Financial Mathematics, 2(1):287–316, ???? 2011. CODEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p287_s1.

Chazal:2018:OPO

[CLP18] M. Chazal, R. Loeffen, and P. Patie. Option pricing in a onedimensional affine term structure model via spectral representations. SIAM Journal on Financial Mathematics, 9(2):634– 664, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Chen:2018:ODS

[CLZ18] Shumin Chen, Zhongfei Li, and Yan Zeng. Optimal dividend strategy for a general diffusion process with time-inconsistent preferences and ruin penalty. SIAM Journal on Financial Mathematics, 9(1):274–314, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Carr:2010:LVE

[CM10] Peter Carr and Dilip B. Madan.

Local volatility enhanced by a jump to default. SIAM Journal on Financial Mathematics, 1 (1):2–15, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Crisan:2012:SBS

[CM12] D. Crisan and K. Manolarakis.

Solving backward stochastic differential equations using the cubature method: Application to nonlinear pricing. SIAM Journal on Financial Mathematics, 3(1): 534–571, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Campolieti:2017:SDM

[CM17] Giuseppe Campolieti and Roman N. Makarov. Solvable diffusion models with linear and mean-reverting nonlinear drifts. SIAM Journal on Financial Mathematics, 8(1):146–170, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Cont:2021:SPD

[CM21] Rama Cont and Marvin S. Müller. A stochastic partial differential equation model for limit order book dynamics. SIAM Journal on Financial Mathematics, 12(2):744– 787, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Czichowsky:2014:TCS

[CMKS14] Christoph Czichowsky, Johannes Muhle-Karbe, and Walter Schachermayer. Transaction costs, shadow prices, and duality in discrete time. SIAM Journal on Financial Mathematics, 5(1): 258–277, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Carmona:2017:SIV

[CMN17] Rene Carmona, Yi Ma, and Sergey Nadtochiy. Simulation of implied volatility surfaces via tangent Lévy models. SIAM Journal on Financial Mathematics, 8(1):171–213, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Cipriano:2021:OPH

[CMP21] Fernanda Cipriano, Nuno F. M. Martins, and Diogo Pereira. Optimal portfolio for the αhypergeometric stochastic volatility model. SIAM Journal on Financial Mathematics, 12(1): 226–253, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Cozma:2018:CES

[CMR18] Andrei Cozma, Matthieu Mariapragassam, and Christoph Reisinger. Convergence of an Euler scheme for a hybrid stochastic-local volatility model with stochastic rates in foreign exchange markets. SIAM Journal on Financial Mathematics, 9(1):127–170, ???? 2018. CO-DEN SJFMBJ. ISSN 1945-497X.

Cozma:2019:CHL

[CMR19] Andrei Cozma, Matthieu Mariapragassam, and Christoph Reisinger. Calibration of a hybrid local-stochastic volatility stochastic rates model with a control variate particle method. SIAM Journal on Financial Mathematics, 10(1):181–213, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Carr:2011:SHU

[CN11] Peter Carr and Sergey Nadtochiy. Static hedging under time-homogeneous diffusions. SIAM Journal on Financial Mathematics, 2(1): 794–838, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p794_s1.

Cheridito:2012:PCS

[CNP12] Patrick Cheridito, Ashkan Nikeghbali, and Eckhard Platen. Processes of class sigma, last passage times, and drawdowns. SIAM Journal on Financial Mathematics, 3(1):280–303, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Cox:2011:RHD

[CO11] A. M. G. Cox and Jan Obloj. Robust hedging of double touch barrier options. SIAM Journal on Financial Mathematics, 2(1):141–182, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL http://epubs.
                                        siam.org/sifin/resource/1/
                                        sjfmbj/v2/i1/p141_s1.

Cotton:2021:IRA

[Cot21] Peter Cotton. Inferring relative ability from winning probability in multientrant contests. SIAM Journal on Financial Mathematics, 12(1):295–317, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Carassus:2019:RSP

[COW19] Laurence Carassus, Jan Obłój, and Johannes Wiesel. The robust superreplication problem: a dynamic approach. SIAM Journal on Financial Mathematics, 10(4):907–941, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Carassus:2022:ERS

[COW22] Laurence Carassus, Jan Obłój, and Johannes Wiesel. Erratum: The robust superreplication problem: a dynamic approach. SIAM Journal on Financial Mathematics, 13(2):653– 655, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1447040.

Chevalier:2022:AOV

[CPZ22] Etienne Chevalier, Sergio Pulido, and Elizabeth Zúñiga. American options in the Volterra Heston model. SIAM Journal on Financial Mathematics, 13(2): 426–458, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M140674X.

Carmona:2010:MEC

[CS10] René Carmona and Ronnie Sircar. Message from the Editorsin-Chief. SIAM Journal on Financial Mathematics, 1(1):1, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Cartea:2021:SPL

[CSB21] Álvaro Cartea and Leandro Sánchez-Betancourt. The shadow price of latency: Improving intraday fill ratios in foreign exchange markets. SIAM Journal on Financial Mathematics, 12(1):254–294, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Chan:2015:FMF

[CSS15] Patrick Chan, Ronnie Sircar, and Michael V. Stein. A feedback model for the financialization of commodity markets. SIAM Journal on Financial Mathematics, 6(1):870–899, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Crepey:2020:WCF

[CSS20] Stéphane Crépey, Wissal Sabbagh, and Shiqi Song. When capital is a funding source: The anticipated backward stochastic differential equations of X-value adjustments. SIAM Journal on Financial Mathematics, 11(1): 99–130, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Chau:2015:MMO

[CT15] Huy N. Chau and Peter Tankov. Market models with optimal arbitrage. SIAM Journal on Financial Mathematics, 6(1):66– 85, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Chevalier:2013:ODI

[CVS13] Etienne Chevalier, Vathana Ly Vath, and Simone Scotti. An optimal dividend and investment control problem under debt constraints. SIAM Journal on Financial Mathematics, 4(1):297– 326, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Cheridito:2012:PHA

[CW12] Patrick Cheridito and Alexander Wugalter. Pricing and hedging in affine models with possibility of default. SIAM Journal on Financial Mathematics, 3(1): 328–350, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Campbell:2022:FPO

[CW22a] Steven Campbell and Ting-Kam Leonard Wong. Functional portfolio optimization in stochastic portfolio theory. SIAM Journal on Financial Mathematics, 13(2):576– 618, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1417715.

Choi:2022:ENT

[CW22b] Jin Hyuk Choi and Kim Weston. Endogenous noise trackers in a radner equilibrium. SIAM Journal on Financial Mathematics, 13(4):1326– 1343, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/22M1483384.

Cohen:2021:ODP

[CY21] Asaf Cohen and Virginia R.

Young. Optimal dividend problem: Asymptotic analysis. SIAM Journal on Financial Mathematics, 12(1):29–46, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Dastgerdi:2020:SPF

[DB20] Maryam Vahid Dastgerdi and Ali Foroush Bastani. Solving parametric fractional differential equations arising from the rough Heston model using quasilinearization and spectral collocation. SIAM Journal on Financial Mathematics, 11(4):1063– 1097, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

DeMarco:2021:HMR

[De 21] Stefano De Marco. On the harmonic mean representation of the implied volatility. SIAM Journal on Financial Mathematics, 12(2):551–565, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

DeMarco:2018:LVC

[DF18] Stefano De Marco and Peter K. Friz. Local volatility, conditioned diffusions, and Varadhan’s formula. SIAM Journal on Financial Mathematics, 9(2): 835–874, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Doldi:2021:CSR

[DF21] Alessandro Doldi and Marco Frittelli. Conditional systemic risk measures. SIAM Journal on Financial Mathematics, 12(4): 1459–1507, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

DeMarco:2017:SIV

[DHJ17] S. De Marco, C. Hillairet, and A. Jacquier. Shapes of implied volatility with positive mass at zero. SIAM Journal on Financial Mathematics, 8(1):709–737, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

DeMarco:2015:LVV

[DHL15] Stefano De Marco and Pierre Henry-Labordère. Linking vanillas and VIX options: a constrained martingale optimal transport problem. SIAM Journal on Financial Mathematics, 6 (1):1171–1194, ???? 2015. CO-DEN SJFMBJ. ISSN 1945-497X.

DelMoral:2011:RSE

[DHOR11] Pierre Del Moral, Peng Hu, Nadia Oudjane, and Bruno Rémillard. On the robustness of the Snell envelope. SIAM Journal on Financial Mathematics, 2(1):587–626, ???? 2011.

CODEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p587_s1.

Duc:2023:HRP

[DJ23] Luu H. Duc and Jürgen Jost.

How rough path lifts affect expected return and volatility: a rough model under transaction cost. SIAM Journal on Financial Mathematics, 14(3): 879–909, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/20M1358670.

Detemple:2018:AOD

[DK18] Jerome Detemple and Yerkin Kitapbayev. American options with discontinuous twolevel caps. SIAM Journal on Financial Mathematics, 9(1):219– 250, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Davis:2011:JDR

[DL11a] Mark Davis and Sébastien Lleo.

Jump-diffusion risk-sensitive asset management I: Diffusion factor model. SIAM Journal on Financial Mathematics, 2 (1):22–54, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p22_s1.

Dia:2011:CCB

[DL11b] El Hadj Aly Dia and Damien Lamberton. Continuity correction for barrier options in jumpdiffusion models. SIAM Journal on Financial Mathematics, 2(1):866–900, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL http://epubs.
                                        siam.org/sifin/resource/1/
                                        sjfmbj/v2/i1/p866_s1.

Dassios:2013:POP

[DL13] Angelos Dassios and Jia Wei Lim. Parisian option pricing: a recursive solution for the density of the Parisian stopping time. SIAM Journal on Financial Mathematics, 4(1):599–615, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Dolinsky:2022:SCU

[DM22] Yan Dolinsky and Shir Moshe. Short communication: Utility indifference pricing with high risk aversion and small linear price impact. SIAM Journal on Financial Mathematics, 13(1): SC12–SC25, ???? 2022. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL https://epubs.
                                        siam.org/doi/10.1137/21M1456431.

Detering:2019:MDC

[DMBPR19] Nils Detering, Thilo Meyer-Brandis, Konstantinos Panagiotou, and Daniel Ritter. Managing default contagion in inhomogeneous financial networks. SIAM Journal on Financial Mathematics, 10(2):578–614, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Detering:2022:SFS

[DMBPR22] Nils Detering, Thilo Meyer-Brandis, Konstantinos Panagiotou, and Daniel Ritter. Suffocating fire sales. SIAM Journal on Financial Mathematics, 13 (1):70–108, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/20M1379800.

Dam:2020:RMI

[DMSS20] Henrik T. Dam, Andrea Macrina, David Skovmand, and David Sloth. Rational models for inflation-linked derivatives. SIAM Journal on Financial Mathematics, 11(4):974– 1006, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Dixon:2020:SCD

[DP20] Matthew Dixon and Nick Polson. Short communication: Deep fundamental factor models. SIAM Journal on Financial Mathematics, 11(3):SC–26– SC–37, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Dumitrescu:2017:GOI

[DQS17] Roxana Dumitrescu, Marie-Claire Quenez, and Agnès Sulem. Game options in an imperfect market with default. SIAM Journal on Financial Mathematics, 8(1):532–559, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Dentcheva:2018:TCR

[DR18] Darinka Dentcheva and Andrzej Ruszczyński. Time-coherent risk measures for continuoustime Markov chains. SIAM Journal on Financial Mathematics, 9(2):690–715, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

dosReis:2022:FUM

[dRP22] Gonçalo dos Reis and Vadim Platonov. Forward utility and market adjustments in relative investment-consumption games of many players. SIAM Journal on Financial Mathematics, 13(3):844–876, ???? 2022. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL https://epubs.
                                        siam.org/doi/10.1137/20M138421X.

Dorsek:2013:ESC

[DT13] Philipp Dörsek and Josef Teichmann. Efficient simulation and calibration of general HJM models by splitting schemes. SIAM Journal on Financial Mathematics, 4(1):575–598, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Dubois:2015:ODP

[DV15] Mathieu S. Dubois and Luitgard A. M. Veraart. Optimal diversification in the presence of parameter uncertainty for a risk averse investor. SIAM Journal on Financial Mathematics, 6(1): 201–241, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Dmitrasinovic-Vidovic:2011:OPM

[DVW11] Gordana Dmitrasinović-Vidović and Antony Ware. Optimal portfolios of mean-reverting instruments. SIAM Journal on Financial Mathematics, 2(1): 748–767, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL http://epubs.
                                        siam.org/sifin/resource/1/
                                        sjfmbj/v2/i1/p748_s1.

Dai:2010:CTM

[DXZ10] Min Dai, Zuo Quan Xu, and Xun Yu Zhou. Continuous-time Markowitz’s model with transaction costs. SIAM Journal on Financial Mathematics, 1(1): 96–125, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Dolinsky:2023:SCE

[DZ23] Yan Dolinsky and Or Zuk.

Short communication: Exponential utility maximization in a discrete time Gaussian framework. SIAM Journal on Financial Mathematics, 14(3):SC31– SC41, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/23M1576074.

Dai:2010:TFT

[DZZ10] M. Dai, Q. Zhang, and Q. J.

Zhu. Trend following trading under a regime switching model. SIAM Journal on Financial Mathematics, 1(1):780–810, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Elizalde:2022:SCC

[EE22] Mauricio Elizalde and Carlos Escudero. Short communication: Chances for the honest in honest versus insider trading. SIAM Journal on Financial Mathematics, 13(2):SC39– SC52, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1439547.

ElEuch:2019:STM

[EFGR19] Omar El Euch, Masaaki Fukasawa, Jim Gatheral, and Mathieu Rosenbaum. Short-term at-the-money asymptotics under stochastic volatility models. SIAM Journal on Financial Mathematics, 10(2):491– 511, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Errais:2010:APP

[EGG10] Eymen Errais, Kay Giesecke, and Lisa R. Goldberg. Affine point processes and portfolio credit risk. SIAM Journal on Financial Mathematics, 1(1):642– 665, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Eckstein:2021:RPH

[EGLO21] Stephan Eckstein, Gaoyue Guo, Tongseok Lim, and Jan Obłój. Robust pricing and hedging of options on multiple assets and its numerics. SIAM Journal on Financial Mathematics, 12(1): 158–188, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Eberlein:2013:DTM

[EGS13] Ernst Eberlein, Zorana Grbac, and Thorsten Schmidt. Discrete tenor models for credit risky portfolios driven by timeinhomogeneous Lévy processes. SIAM Journal on Financial Mathematics, 4(1):616–649, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

ElKaroui:2015:DAM

[EJJ15] Nicole El Karoui, Monique Jeanblanc, and Ying Jiao. Density approach in modeling successive defaults. SIAM Journal on Financial Mathematics, 6(1):1–21, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

ElAmrani:2021:SCD

[EJM21] Mehdi El Amrani, Antoine Jacquier, and Claude Martini. Short communication: Dynamics of symmetric SSVI smiles and implied volatility bubbles. SIAM Journal on Financial Mathematics, 12(2):SC1–SC15, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

ElKaroui:2013:RRH

[El 13] Noureddine El Karoui. On the realized risk of high-dimensional Markowitz portfolios. SIAM Journal on Financial Mathematics, 4(1):737–783, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

ElKaroui:2021:RDU

[EM21] Nicole El Karoui and Mohamed Mrad. Recover dynamic utility from observable process: Application to the economic equilibrium. SIAM Journal on Financial Mathematics, 12(1):189– 225, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Ekeland:2012:TCP

[EMP12] Ivar Ekeland, Oumar Mbodji, and Traian A. Pirvu. Timeconsistent portfolio management. SIAM Journal on Financial Mathematics, 3(1): 1–32, ???? 2012. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL http://epubs.
                                        siam.org/sifin/resource/1/
                                        sjfmbj/v3/i1/p1_s1.

Elliott:2021:FRD

[EMW21] Robert J. Elliott, Dilip B. Madan, and King Wang. Filtering response directions. SIAM Journal on Financial Mathematics, 12(3):1285–1306, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Ekstrom:2016:OLA

[EV16] Erik Ekström and Juozas Vaicenavicius. Optimal liquidation of an asset under drift uncertainty. SIAM Journal on Financial Mathematics, 7(1):357–381, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Feinstein:2019:OPD

[Fei19] Zachary Feinstein. Obligations with physical delivery in a multilayered financial network. SIAM Journal on Financial Mathematics, 10(4):877–906, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Feinstein:2022:SCC

[Fei22] Zachary Feinstein. Short communication: Clearing prices under margin calls and the short squeeze. SIAM Journal on Financial Mathematics, 13(4): SC113–SC122, ???? 2022. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL https://epubs.
                                        siam.org/doi/10.1137/22M147877X.

Feng:2010:SMA

[FFF10] Jin Feng, Martin Forde, and Jean-Pierre Fouque. Shortmaturity asymptotics for a fast mean-reverting Heston stochastic volatility model. SIAM Journal on Financial Mathematics, 1 (1):126–141, ???? 2010. CO-DEN SJFMBJ. ISSN 1945-497X.

Fouque:2018:OPU

[FH18] Jean-Pierre Fouque and Ruimeng Hu. Optimal portfolio under fast mean-reverting fractional stochastic environment. SIAM Journal on Financial Mathematics, 9(2):564–601, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Feinstein:2023:CCO

[FH23] Zachary Feinstein and Thomas R.

Hurd. Contingent convertible obligations and financial stability. SIAM Journal on Financial Mathematics, 14(1):158– 187, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/22M1498954.

Fouque:2022:SSA

[FHS22] Jean-Pierre Fouque, Ruimeng Hu, and Ronnie Sircar. Suband supersolution approach to accuracy analysis of portfolio optimization asymptotics in multiscale stochastic factor markets. SIAM Journal on Financial Mathematics, 13(1):109– 128, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1428625.

Fouque:2013:SMI

[FI13] Jean-Pierre Fouque and Tomoyuki Ichiba. Stability in a model of interbank lending. SIAM Journal on Financial Mathematics, 4(1):784–803, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Farahany:2020:MLP

[FJJ20] David Farahany, Kenneth R. Jackson, and Sebastian Jaimungal. Mixing LSMC and PDE methods to price Bermudan options. SIAM Journal on Financial Mathematics, 11(1):201– 239, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Fouque:2011:SDO

[FJL11] Jean-Pierre Fouque, Sebastian Jaimungal, and Matthew J. Lorig. Spectral decomposition of option prices in fast meanreverting stochastic volatility models. SIAM Journal on Financial Mathematics, 2(1): 665–691, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL http://epubs.
                                        siam.org/sifin/resource/1/
                                        sjfmbj/v2/i1/p665_s1.

Forde:2012:STS

[FJL12] Martin Forde, Antoine Jacquier, and Roger Lee. The small-time smile and term structure of implied volatility under the Heston model. SIAM Journal on Financial Mathematics, 3(1):690–708, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Fontanela:2021:SCQ

[FJO21] Filipe Fontanela, Antoine Jacquier, and Mugad Oumgari. Short communication: a quantum algorithm for linear PDEs arising in finance. SIAM Journal on Financial Mathematics, 12(4):SC98–SC114, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Fouque:2022:OTS

[FJS22] Jean-Pierre Fouque, Sebastian Jaimungal, and Yuri F. Saporito. Optimal trading with signals and stochastic price impact. SIAM Journal on Financial Mathematics, 13(3):944– 968, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1394473.

Fuh:2021:CRP

[FK21] Cheng-Der Fuh and Chu-Lan Michael Kao. Credit risk propagation in structural-form models. SIAM Journal on Financial Mathematics, 12(4): 1340–1373, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Filipovic:2012:ACR

[FKV12] Damir Filipović, Michael Kupper, and Nicolas Vogelpoth. Approaches to conditional risk. SIAM Journal on Financial Mathematics, 3(1):402–432, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Fouque:2011:FMR

[FL11] Jean-Pierre Fouque and Matthew J. Lorig. A fast mean-reverting correction to Heston’s stochastic volatility model. SIAM Journal on Financial Mathematics, 2(1):221–254, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL http://epubs.
                                        siam.org/sifin/resource/1/
                                        sjfmbj/v2/i1/p221_s1.

Feng:2013:IAC

[FL13a] Liming Feng and Xiong Lin. Inverting analytic characteristic functions and financial applications. SIAM Journal on Financial Mathematics, 4(1):372–398, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Feng:2013:PBO

[FL13b] Liming Feng and Xiong Lin. Pricing Bermudan options in Lévy process models. SIAM Journal on Financial Mathematics, 4(1):474–493, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Figueroa-Lopez:2012:SMS

[FLF12] José E. Figueroa-López and Martin Forde. The smallmaturity smile for exponential Lévy models. SIAM Journal on Financial Mathematics, 3 (1):33–65, ???? 2012. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL http://epubs.
                                        siam.org/sifin/resource/1/
                                        sjfmbj/v3/i1/p33_s1.

Figueroa-Lopez:2018:STE

[FLGL18] José E. Figueroa-López, Ruoting Gong, and Matthew Lorig. Short-time expansions for call options on leveraged ETFs under exponential Lévy models with local volatility. SIAM Journal on Financial Mathematics, 9 (1):347–380, ???? 2018. CO-DEN SJFMBJ. ISSN 1945-497X.

Fadina:2023:OAR

[FLW23] Tolulope Fadina, Peng Liu, and Ruodu Wang. One axiom to rule them all: a minimalist axiomatization of quantiles. SIAM Journal on Financial Mathematics, 14(2):644– 662, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/22M1531567.

Frittelli:2011:DRQ

[FM11] Marco Frittelli and Marco Maggis. Dual representation of quasi-convex conditional maps. SIAM Journal on Financial Mathematics, 2 (1):357–382, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p357_s1.

Faidi:2011:MRU

[FMM11a] Wahid Faidi, Anis Matoussi, and Mohamed Mnif. Maximization of recursive utilities: a dynamic maximum principle approach. SIAM Journal on Financial Mathematics, 2(1):

1014–1041, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL http://epubs.
                                        siam.org/sifin/resource/1/
                                        sjfmbj/v2/i1/p1014_s1.

Fusai:2011:PDM

[FMM11b] Gianluca Fusai, Daniele Marazzina, and Marina Marena. Pricing discretely monitored Asian options by maturity randomization. SIAM Journal on Financial Mathematics, 2(1): 383–403, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL http://epubs.
                                        siam.org/sifin/resource/1/
                                        sjfmbj/v2/i1/p383_s1.

Fouque:2018:UVM

[FN18] Jean-Pierre Fouque and Ning Ning. Uncertain volatility models with stochastic bounds. SIAM Journal on Financial Mathematics, 9(4):1175–1207, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Fang:2011:FBV

[FO11] Fang Fang and Cornelis W. Oosterlee. A Fourier-based valuation method for Bermudan and barrier options under Heston’s model. SIAM Journal on Financial Mathematics, 2 (1):439–463, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL http://epubs.
                                        siam.org/sifin/resource/1/
                                        sjfmbj/v2/i1/p439_s1.

Fox:2021:BPG

[FÖ21] Jamie Fox and Giray Ökten.

Brownian path generation and polynomial chaos. SIAM Journal on Financial Mathematics, 12(2):724–743, ???? 2021. CO-DEN SJFMBJ. ISSN 1945-497X.

Fontana:2023:SCC

[Fon23] Claudio Fontana. Short communication: Caplet pricing in affine models for alternative riskfree rates. SIAM Journal on Financial Mathematics, 14(1): SC1–SC16, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/22M1513691.

Forsyth:2020:MMC

[For20] Peter A. Forsyth. Multiperiod mean conditional value at risk asset allocation: Is it advantageous to be time consistent? SIAM Journal on Financial Mathematics, 11(2):358– 384, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Fodra:2015:HFT

[FP15] Pietro Fodra and Huyên Pham.

High frequency trading and asymptotics for small risk aversion in a Markov renewal model. SIAM Journal on Financial Mathematics, 6(1):656–684, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Feinstein:2018:SEN

[FPR^{+}18] Zachary Feinstein, Weijie Pang, Birgit Rudloff, Eric Schaanning, Stephan Sturm, and Mackenzie Wildman. Sensitivity of the Eisenberg–Noe clearing vector to individual interbank liabilities. SIAM Journal on Financial Mathematics, 9(4):1286– 1325, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Fouque:2014:AOP

[FR14] Jean-Pierre Fouque and Bin Ren. Approximation for option prices under uncertain volatility. SIAM Journal on Financial Mathematics, 5(1):360–383, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Frikha:2014:SRM

[Fri14] N. Frikha. Shortfall risk minimization in discrete time financial market models. SIAM Journal on Financial Mathematics, 5 (1):384–414, ???? 2014. CO-DEN SJFMBJ. ISSN 1945-497X.

Feinstein:2017:MSR

[FRW17] Zachary Feinstein, Birgit Rudloff, and Stefan Weber. Measures of systemic risk. SIAM Journal on Financial Mathematics, 8(1): 672–708, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Feinstein:2021:SCD

[FS21] Zachary Feinstein and Andreas Søjmark. Short communication: Dynamic default contagion in heterogeneous interbank systems. SIAM Journal on Financial Mathematics, 12(4):SC83– SC97, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Forzani:2015:LSC

[FT15] Liliana Forzani and Carlos F.

Tolmasky. On the level-slope-curvature effect in yield curves and eventual total positivity. SIAM Journal on Financial Mathematics, 6(1):900–918, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Fries:2020:AVF

[FT20] Christian Fries and Lorenzo Torricelli. An analytical valuation framework for financial assets with trading suspensions. SIAM Journal on Financial Mathematics, 11(2):566–592, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Fujii:2022:SCM

[FT22] Masaaki Fujii and Akihiko Takahashi. Strong convergence to the mean field limit of a finite agent equilibrium. SIAM Journal on Financial Mathematics, 13(2): 459–490, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1441055.

Filipovic:2010:TSM

[FTT10] Damir Filipović, Stefan Tappe, and Josef Teichmann. Term structure models driven by Wiener processes and Poisson measures: Existence and positivity. SIAM Journal on Financial Mathematics, 1(1):523–554, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Filipovic:2018:EST

[FW18] Damir Filipović and Sander Willems. Exact smooth termstructure estimation. SIAM Journal on Financial Mathematics, 9(3):907–929, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Forde:2016:STA

[FZ16] Martin Forde and Hongzhong Zhang. Small-time asymptotics for basket options — the bivariate SABR model and the hyperbolic heat kernel on H^{3}. SIAM Journal on Financial Mathematics, 7(1):448–476, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Forde:2017:ARS

[FZ17] Martin Forde and Hongzhong Zhang. Asymptotics for rough stochastic volatility models. SIAM Journal on Financial Mathematics, 8(1):114–145, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Feng:2023:CMS

[FZ23] Qi Feng and Jianfeng Zhang. Cubature method for stochastic Volterra integral equations. SIAM Journal on Financial Mathematics, 14(4):959– 1003, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M146889X.

Gassiat:2023:WER

[Gas23] Paul Gassiat. Weak error rates of numerical schemes for rough volatility. SIAM Journal on Financial Mathematics, 14(2): 475–496, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/22M1485760.

Gnoatto:2014:AMM

[GG14] Alessandro Gnoatto and Martino Grasselli. An affine multicurrency model with stochastic volatility and stochastic interest rates. SIAM Journal on Financial Mathematics, 5(1):493–531, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Gass:2018:FGS

[GG18] Maximilian Gaß and Kathrin Glau. A flexible Galerkin scheme for option pricing in Lévy models. SIAM Journal on Financial Mathematics, 9(3):930–965, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Gass:2017:MPF

[GGM17] Maximilian Gaß, Kathrin Glau, and Maximilian Mair. Magic points in finance: Empirical integration for parametric option pricing. SIAM Journal on Financial Mathematics, 8(1):766– 803, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Gomes:2023:RSM

[GGR23] Diogo Gomes, Julian Gutierrez, and Ricardo Ribeiro. A random-supply mean field game price model. SIAM Journal on Financial Mathematics, 14(1): 188–222, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1443923.

Guasoni:2021:SCA

[GHK21] Paolo Guasoni, Yu-Jui Huang, and Saeed Khalili. Short communication: American student loans: Repayment and valuation. SIAM Journal on Financial Mathematics, 12(2):SC16– SC30, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Guo:2016:GAF

[GJMN16] Gaoyue Guo, Antoine Jacquier, Claude Martini, and Leo Neufcourt. Generalized arbitrage-free SVI volatility surfaces. SIAM Journal on Financial Mathematics, 7(1):619–641, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Guennoun:2018:ABF

[GJRS18] Hamza Guennoun, Antoine Jacquier, Patrick Roome, and Fangwei Shi. Asymptotic behavior of the fractional Heston model. SIAM Journal on Financial Mathematics, 9(3):1017– 1045, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Goodman:2010:CFR

[GK10] Victor Goodman and Kyounghee Kim. Common forward rate volatility. SIAM Journal on Financial Mathematics, 1(1): 212–229, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Garreau:2016:SJT

[GK16] Pierre Garreau and Alec Kercheval. A structural jump threshold framework for credit risk. SIAM Journal on Financial Mathematics, 7(1):642–673, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Gurdogan:2022:MAP

[GK22] Hubeyb Gurdogan and Alec Kercheval. Multiple anchor point shrinkage for the sample covariance matrix. SIAM Journal on Financial Mathematics, 13(3):1112–1143, ???? 2022. CODEN SJFMBJ. ISSN 1945-

497X. URL https://epubs.
siam.org/doi/10.1137/21M1446411.

Giesecke:2010:EES

[GKMT10] K. Giesecke, H. Kakavand, M. Mousavi, and H. Takada. Exact and efficient simulation of correlated defaults. SIAM Journal on Financial Mathematics, 1(1):868–896, ???? 2010. CODEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v1/i1/p868_s1.

Griessler:2014:COD

[GKR14] Claus Griessler and Martin Keller-Ressel. Convex order of discrete, continuous, and predictable quadratic variation and applications to options on variance. SIAM Journal on Financial Mathematics, 5(1):1–19, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Glau:2020:LRT

[GKS20] Kathrin Glau, Daniel Kressner, and Francesco Statti. Low-rank tensor approximation for Chebyshev interpolation in parametric option pricing. SIAM Journal on Financial Mathematics, 11(3): 897–927, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Gramacy:2015:SDO

[GL15] Robert B. Gramacy and Michael Ludkovski. Sequential design for optimal stopping problems. SIAM Journal on Financial Mathematics, 6(1):748–775, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Gapeev:2022:PAS

[GL22] Pavel V. Gapeev and Libo Li. Perpetual American standard and lookback options with event risk and asymmetric information. SIAM Journal on Financial Mathematics, 13(3):773– 801, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1396848.

Gueant:2012:OPL

[GLFT12] Olivier Guéant, Charles-Albert Lehalle, and Joaquin Fernandez-Tapia. Optimal portfolio liquidation with limit orders. SIAM Journal on Financial Mathematics, 3(1):740–764, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Guo:2022:JMC

[GLOW22] Ivan Guo, Grégoire Loeper, Jan Obłój, and Shiyi Wang. Joint modeling and calibration of SPX and VIX by optimal transport. SIAM Journal on Financial Mathematics, 13(1): 1–31, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/20M1375905.

Guan:2020:OIP

[GLZ20] Chonghu Guan, Xun Li, and Wenxin Zhou. An optimal investment problem with nonsmooth and nonconcave utility over a finite time horizon. SIAM Journal on Financial Mathematics, 11(2):411–436, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

GarciadelMolino:2020:MKM

[GMBB20] Luis Carlos Garcia del Molino, Iacopo Mastromatteo, Michael Benzaquen, and Jean-Philippe Bouchaud. The multivariate Kyle model: More is different. SIAM Journal on Financial Mathematics, 11(2):327– 357, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Guasoni:2019:TFB

[GNR19] Paolo Guasoni, Zsolt Nika, and Miklós Rásonyi. Trading fractional Brownian motion. SIAM Journal on Financial Mathematics, 10(3):769–789, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Goodman:2011:ORT

[GO11a] Jonathan Goodman and Daniel N.

Ostrov. An option to reduce transaction costs. SIAM Journal on Financial Mathematics, 2(1):512–537, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL http://epubs.
                                        siam.org/sifin/resource/1/
                                        sjfmbj/v2/i1/p512_s1.

Grzelak:2011:HMS

[GO11b] Lech A. Grzelak and Cornelis W. Oosterlee. On the Heston model with stochastic interest rates. SIAM Journal on Financial Mathematics, 2(1): 255–286, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL http://epubs.
                                        siam.org/sifin/resource/1/
                                        sjfmbj/v2/i1/p255_s1.

Gobet:2015:AAB

[GP15] Emmanuel Gobet and Stefano Pagliarani. Analytical approximations of BSDEs with nonsmooth driver. SIAM Journal on Financial Mathematics, 6(1): 919–958, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Goldberg:2022:DB

[GPS22] Lisa R. Goldberg, Alex Papanicolaou, and Alex Shkolnik. The dispersion bias. SIAM Journal on Financial Mathematics, 13 (2):521–550, ???? 2022. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL https://epubs.
                                        siam.org/doi/10.1137/21M144058X.

Grbac:2015:ALM

[GPSS15] Zorana Grbac, Antonis Papapantoleon, John Schoenmakers, and David Skovmand. Affine LIBOR models with multiple curves: Theory, examples and calibration. SIAM Journal on Financial Mathematics, 6(1):

984–1025, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Garnier:2013:LDM

[GPY13] Josselin Garnier, George Papanicolaou, and Tzu-Wei Yang. Large deviations for a mean field model of systemic risk. SIAM Journal on Financial Mathematics, 4(1):151–184, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Grigorova:2020:EON

[GQS20] Miryana Grigorova, Marie-Claire Quenez, and Agnès Sulem. European options in a nonlinear incomplete market model with default. SIAM Journal on Financial Mathematics, 11(3):849–880, ???? 2020. CO-DEN SJFMBJ. ISSN 1945-497X.

Giles:2012:SFD

[GR12] Michael B. Giles and Christoph Reisinger. Stochastic finite differences and multilevel Monte Carlo for a class of SPDEs in finance. SIAM Journal on Financial Mathematics, 3(1):572–592, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Gueant:2014:VEG

[GR14] Olivier Guéant and Guillaume Royer. VWAP execution and guaranteed VWAP. SIAM Journal on Financial Mathematics, 5 (1):445–471, ???? 2014. CO-DEN SJFMBJ. ISSN 1945-497X.

Glasserman:2012:QTA

[GS12] Paul Glasserman and Sira Suchintabandid. Quadratic transform approximation for CDO pricing in multifactor models. SIAM Journal on Financial Mathematics, 3(1):137– 162, ???? 2012. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL http://epubs.
                                        siam.org/sifin/resource/1/
                                        sjfmbj/v3/i1/p137_s1.

Garnier:2017:CBS

[GS17] Josselin Garnier and Knut Sølna. Correction to Black– Scholes formula due to fractional stochastic volatility. SIAM Journal on Financial Mathematics, 8(1):560–588, ???? 2017. CO-DEN SJFMBJ. ISSN 1945-497X.

Garnier:2020:OHU

[GS20] Josselin Garnier and Knut Sølna. Optimal hedging under fast-varying stochastic volatility. SIAM Journal on Financial Mathematics, 11(1):274– 325, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Gnoatto:2021:CCV

[GS21] Alessandro Gnoatto and Nicole Seiffert. Cross currency valuation and hedging in the multiple curve framework. SIAM Journal on Financial Mathematics, 12(3):967–1012, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Guasoni:2019:SCI

[GTW19] Paolo Guasoni, Antonella Tolomeo, and Gu Wang. Should commodity investors follow commodities’ prices? SIAM Journal on Financial Mathematics, 10(2): 466–490, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Gulisashvili:2010:AFE

[Gul10] Archil Gulisashvili. Asymptotic formulas with error estimates for call pricing functions and the implied volatility at extreme strikes. SIAM Journal on Financial Mathematics, 1(1):609–641, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Gulisashvili:2018:LDP

[Gul18] Archil Gulisashvili. Large deviation principle for Volterra type fractional stochastic volatility models. SIAM Journal on Financial Mathematics, 9(3):1102– 1136, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Guyon:2022:VFB

[Guy22] Julien Guyon. The VIX future in Bergomi models: Fast approximation formulas and joint calibration with S&P 500 skew. SIAM Journal on Financial Mathematics, 13(4):1418– 1485, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1437408.

Gulisashvili:2015:AAS

[GV15] Archil Gulisashvili and Josep Vives. Asymptotic analysis of stock price densities and implied volatilities in mixed stochastic models. SIAM Journal on Financial Mathematics, 6(1):158– 188, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Granelli:2016:MVR

[GV16] Andrea Granelli and Almut E. D. Veraart. Modeling the variance risk premium of equity indices: The role of dependence and contagion. SIAM Journal on Financial Mathematics, 7(1): 382–417, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Guo:2015:OEM

[GZ15] Xin Guo and Mihail Zervos. Optimal execution with multiplicative price impact. SIAM Journal on Financial Mathematics, 6(1): 281–306, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Hepperger:2010:OPH

[Hep10] Peter Hepperger. Option pricing in Hilbert space-valued jumpdiffusion models using partial integro-differential equations. SIAM Journal on Financial Mathematics, 1(1):454–489, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Hinz:2010:SCC

[HF10] Juri Hinz and Max Fehr. Storage costs in commodity option pricing. SIAM Journal on Financial Mathematics, 1(1):729–751, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Hamel:2010:DSV

[HH10] Andreas H. Hamel and Frank Heyde. Duality for set-valued measures of risk. SIAM Journal on Financial Mathematics, 1 (1):66–95, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Horvath:2020:VOR

[HJT20] Blanka Horvath, Antoine Jacquier, and Peter Tankov. Volatility options in rough volatility models. SIAM Journal on Financial Mathematics, 11(2): 437–469, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Hambly:2017:SEE

[HK17a] Ben Hambly and Nikolaos Kolliopoulos. Stochastic evolution equations for large portfolios of stochastic volatility models. SIAM Journal on Financial Mathematics, 8(1):962– 1014, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X. See erratum [HK19].

Horst:2017:WLL

[HK17b] Ulrich Horst and Dörte Kreher.

A weak law of large numbers for a limit order book model with fully state dependent order dynamics. SIAM Journal on Financial Mathematics, 8(1): 314–343, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Hayashi:2018:WBM

[HK18] Takaki Hayashi and Yuta Koike.

Wavelet-based methods for highfrequency lead-lag analysis. SIAM Journal on Financial Mathematics, 9(4):1208–1248, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Hambly:2019:ESE

[HK19] Ben Hambly and Nikolaos Kolliopoulos. Erratum: Stochastic Evolution Equations for Large Portfolios of Stochastic Volatility Models. SIAM Journal on Financial Mathematics, 10(3): 857–876, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X. See [HK17a].

Henderson:2020:ESO

[HKMR20] Vicky Henderson, Kamil Kladívko, Michael Monoyios, and Christoph Reisinger. Executive stock option exercise with full and partial information on a drift change point. SIAM Journal on Financial Mathematics, 11(4):1007– 1062, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

He:2017:RDU

[HKZ17] Xue Dong He, Roy Kouwenberg, and Xun Yu Zhou. Rankdependent utility and risk taking in complete markets. SIAM Journal on Financial Mathematics, 8(1):214–239, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Henry-Labordere:2016:DAS

[HLLR16] Pierre Henry-Labordère, Christian Litterer, and Zhenjie Ren. A dual algorithm for stochastic control problems: Applications to uncertain volatility models and CVA. SIAM Journal on Financial Mathematics, 7(1):159– 182, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Herrmann:2020:IMH

[HMKSY20] Sebastian Herrmann, Johannes Muhle-Karbe, Dapeng Shang, and Chen Yang. Inventory management for high-frequency trading with imperfect competition. SIAM Journal on Financial Mathematics, 11(1):1– 26, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Horst:2014:WCS

[HN14] Ulrich Horst and Felix Naujokat. When to cross the spread? Trading in two-sided limit order books. SIAM Journal on Financial Mathematics, 5(1):278–315, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Huang:2011:SAE

[HO11] Xinzheng Huang and Cornelis W. Oosterlee. Saddlepoint approximations for expectations and an application to CDO pricing. SIAM Journal on Financial Mathematics, 2 (1):692–714, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p692_s1.

Howison:2012:AAA

[How12] Sam Howison. Asymptotic approximations for Asian, European, and American options with discrete averaging or discrete dividend/coupon payments. SIAM Journal on Financial Mathematics, 3(1):215–241, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Huang:2017:EDM

[HR17] Weibing Huang and Mathieu Rosenbaum. Ergodicity and diffusivity of Markovian order book models: a general framework. SIAM Journal on Financial Mathematics, 8(1):874–900, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Horvath:2018:DFF

[HR18] Blanka Horvath and Oleg Reichmann. Dirichlet forms and finite element methods for the SABR model. SIAM Journal on Financial Mathematics, 9(2):716–754, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Howison:2013:ENP

[HRW13] S. D. Howison, C. Reisinger, and J. H. Witte. The effect of nonsmooth payoffs on the penalty approximation of American options. SIAM Journal on Financial Mathematics, 4(1):539–574, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Howison:2012:RNP

[HS12] Sam Howison and Daniel Schwarz. Risk-neutral pricing of financial instruments in emission markets: a structural approach. SIAM Journal on Financial Mathematics, 3(1):709– 739, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Ho:2015:WEN

[HSX15] Michael Ho, Zheng Sun, and Jack Xin. Weighted elastic net penalized mean-variance portfolio design and computation. SIAM Journal on Financial Mathematics, 6(1):1220– 1244, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Hu:2023:CMM

[HSX23] Ying Hu, Xiaomin Shi, and Zuo Quan Xu. Constrained monotone mean-variance problem with random coefficients. SIAM Journal on Financial Mathematics, 14(3):838–854, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X. URL https:/

/epubs.siam.org/doi/10.1137/
22M154418X.

Huang:2017:WCP

[HSZ17] Yao Tung Huang, Qingshuo Song, and Harry Zheng. Weak convergence of path-dependent SDEs in basket credit default swap pricing with contagion risk. SIAM Journal on Financial Mathematics, 8(1):1–27, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Haugh:2014:DPE

[HW14] Martin Haugh and Chun Wang.

Dynamic portfolio execution and information relaxations. SIAM Journal on Financial Mathematics, 5(1):316–359, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Han:2021:TIR

[HW21] Bingyan Han and Hoi Ying Wong. Time-inconsistency with rough volatility. SIAM Journal on Financial Mathematics, 12 (4):1553–1595, ???? 2021. CO-DEN SJFMBJ. ISSN 1945-497X.

Hou:2016:RMM

[HX16] Danlin Hou and Zuo Quan Xu. A robust Markowitz mean-variance portfolio selection model with an intractable claim. SIAM Journal on Financial Mathematics, 7(1):124–151, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Horst:2019:SLL

[HX19] Ulrich Horst and Wei Xu. A scaling limit for limit order books driven by Hawkes processes. SIAM Journal on Financial Mathematics, 10(2):350– 393, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Hurd:2010:FTM

[HZ10] T. R. Hurd and Zhuowei Zhou. A Fourier transform method for spread option pricing. SIAM Journal on Financial Mathematics, 1(1):142–157, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Hobson:2016:OCS

[HZ16] David Hobson and Yeqi Zhu. Optimal consumption and sale strategies for a risk averse agent. SIAM Journal on Financial Mathematics, 7(1):674–719, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Huang:2017:OIG

[HZK17] Yao Tung Huang, Pingping Zeng, and Yue Kuen Kwok. Optimal initiation of guaranteed lifelong withdrawal benefit with dynamic withdrawals. SIAM Journal on Financial Mathematics, 8(1):804–840, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Imkeller:2014:TS

[IR14] Nora Imkeller and L. C. G.

Rogers. Trading to stops. SIAM Journal on Financial Mathematics, 5(1):753–781, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Ishii:2021:EUV

[IR21] Hitoshi Ishii and Alexandre Roch. Existence and uniqueness of viscosity solutions of an integro-differential equation arising in option pricing. SIAM Journal on Financial Mathematics, 12(2):604–640, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Jones:2016:OMP

[JC16] Chris Jones and Xinfu Chen.

Optimal mortgage prepayment under the Cox–Ingersoll–Ross model. SIAM Journal on Financial Mathematics, 7(1):552–566, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Jarrow:2011:HDA

[JKP11] Robert Jarrow, Younes Kchia, and Philip Protter. How to detect an asset bubble. SIAM Journal on Financial Mathematics, 2(1):839–865, ???? 2011. CODEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.

                                        siam.org/sifin/resource/1/
                                        sjfmbj/v2/i1/p839_s1.

Jacquier:2018:IVS

[JKR18] Antoine Jacquier and Martin Keller-Ressel. Implied volatility in strict local martingale models. SIAM Journal on Financial Mathematics, 9(1):171–189, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Jacquier:2013:SCL

[JL13] Antoine Jacquier and Matthew Lorig. The smile of certain Lévytype models. SIAM Journal on Financial Mathematics, 4(1): 804–830, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Jarrow:2015:IEU

[JL15] Robert A. Jarrow and Martin Larsson. Informational efficiency under short sale constraints. SIAM Journal on Financial Mathematics, 6(1):804– 824, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Jacquier:2018:OLL

[JL18] Antoine Jacquier and Hao Liu. Optimal liquidation in a level-I limit order book for large-tick Stocks. SIAM Journal on Financial Mathematics, 9(3):875–906, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Jarrow:2020:IET

[JL20a] Robert Jarrow and Martin Larsson. Informational efficiency with trading constraints: a characterization. SIAM Journal on Financial Mathematics, 11(4):

959–973, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Jeanblanc:2020:CCD

[JL20b] Monique Jeanblanc and Libo Li. Characteristics and constructions of default times. SIAM Journal on Financial Mathematics, 11(3):720–749, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Janecek:2020:OIH

[JLS20] Karel Janecek, Zheng Li, and Mihai Sîrbu. Optimal investment with high-watermark fee in a multidimensional jump diffusion model. SIAM Journal on Financial Mathematics, 11(3): 750–787, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Jaber:2021:MPS

[JMP21] Eduardo Abi Jaber, Enzo Miller, and Huyên Pham. Markowitz portfolio selection for multivariate affine and quadratic Volterra models. SIAM Journal on Financial Mathematics, 12(1):369– 409, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Jacquier:2023:DCD

[JO23] Antoine Jacquier and Mugad Oumgari. Deep curve-dependent PDEs for affine rough volatility. SIAM Journal on Financial Mathematics, 14(2):353– 382, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/19M1267805.

Jarrow:2015:LSH

[JP15] Robert Jarrow and Philip Protter. Liquidity suppliers and high frequency trading. SIAM Journal on Financial Mathematics, 6(1):189–200, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Jaimungal:2022:RRA

[JPWT22] Sebastian Jaimungal, Silvana M. Pesenti, Ye Sheng Wang, and Hariom Tatsat. Robust risk-aware reinforcement learning. SIAM Journal on Financial Mathematics, 13(1):213– 226, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M144640X.

Jia:2019:DPO

[JPZ19] Longjie Jia, Martijn Pistorius, and Harry Zheng. Dynamic portfolio optimization with looping contagion risk. SIAM Journal on Financial Mathematics, 10(1):1–36, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Jacquier:2013:SMH

[JR13] Antoine Jacquier and Patrick Roome. The small-maturity Heston forward smile. SIAM Journal on Financial Mathematics, 4 (1):831–856, ???? 2013. CO-DEN SJFMBJ. ISSN 1945-497X.

Jacquier:2015:AFI

[JR15] Antoine Jacquier and Patrick Roome. Asymptotics of forward implied volatility. SIAM Journal on Financial Mathematics, 6(1): 307–351, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Jaimungal:2011:LBC

[JS11] Sebastian Jaimungal and Vladimir Surkov. Lévy-based crosscommodity models and derivative valuation. SIAM Journal on Financial Mathematics, 2 (1):464–487, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p464_s1.

Jacquier:2019:RHM

[JS19] Antoine Jacquier and Fangwei Shi. The randomized Heston model. SIAM Journal on Financial Mathematics, 10(1):89–129, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Johnson:2011:BBA

[JSDN11] Paul V. Johnson, Nicholas J.

Sharp, Peter W. Duck, and David P. Newton. A bridge between American and European options: The “Ameripean” delayed-exercise model. SIAM Journal on Financial Mathematics, 2(1):965–988, ???? 2011. CODEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p965_s1.

Jena:2011:AOM

[JT11a] Rudra P. Jena and Peter Tankov. Arbitrage opportunities in misspecified stochastic volatility models. SIAM Journal on Financial Mathematics, 2(1):317–341, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL http://epubs.
                                        siam.org/sifin/resource/1/
                                        sjfmbj/v2/i1/p317_s1.

Jordan:2011:AAD

[JT11b] Richard Jordan and Charles Tier. Asymptotic approximations to deterministic and stochastic volatility models. SIAM Journal on Financial Mathematics, 2(1):935–964, ???? 2011. CODEN SJFMBJ. ISSN 1945-497X. URL http://

                                        epubs.siam.org/sifin/resource/
                                        1/sjfmbj/v2/i1/p935_s1.

Jacquier:2020:ADO

[JT20] Antoine Jacquier and Lorenzo Torricelli. Anomalous diffusions in option prices: Connecting trade duration and the volatility term structure. SIAM Journal on Financial Mathematics, 11(4):1137–1167, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Jusselin:2021:OMM

[Jus21] Paul Jusselin. Optimal market making with persistent order flow. SIAM Journal on Financial Mathematics, 12(3): 1150–1200, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Jourdain:2011:REB

[JV11] B. Jourdain and M. H. Vellekoop. Regularity of the exercise boundary for American put options on assets with discrete dividends. SIAM Journal on Financial Mathematics, 2(1): 538–561, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p538_s1.

Kallblad:2020:BII

[Käl20] Sigrid Källblad. Black’s inverse investment problem and forward criteria with consumption. SIAM Journal on Financial Mathematics, 11(2):494– 525, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Kardaras:2015:VPE

[Kar15] Constantinos Kardaras. Valuation and parities for exchange options. SIAM Journal on Financial Mathematics, 6(1):140– 157, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Kohatsu-Higa:2010:WKB

[KHOL10] A. Kohatsu-Higa and S. Ortiz-Latorre. Weak Kyle–back equilibrium models for Max and ArgMax. SIAM Journal on Financial Mathematics, 1(1):179– 211, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Kirkby:2015:EOP

[Kir15] J. Lars Kirkby. Efficient option pricing by frame duality with the Fast Fourier Transform. SIAM Journal on Financial Mathematics, 6(1):713–747, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Kirkby:2016:ETM

[Kir16] J. Lars Kirkby. An efficient transform method for Asian option pricing. SIAM Journal on Financial Mathematics, 7(1): 845–892, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Kalsi:2020:OER

[KLA20] Jasdeep Kalsi, Terry Lyons, and Imanol Perez Arribas. Optimal execution with rough path signatures. SIAM Journal on Financial Mathematics, 11(2): 470–493, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Kreher:2023:JDA

[KM23] Dörte Kreher and Cassandra Milbradt. Jump diffusion approximation for the price dynamics of a fully state dependent limit order book model. SIAM Journal on Financial Mathematics, 14(1):1– 51, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/20M1380922.

Kumar:2018:AAO

[KN18] Rohini Kumar and Hussein Nasralah. Asymptotic approximation of optimal portfolio for small time horizons. SIAM Journal on Financial Mathematics, 9 (2):755–774, ???? 2018. CO-DEN SJFMBJ. ISSN 1945-497X.

Kolkiewicz:2015:SDH

[Kol15] Adam W. Kolkiewicz. On suboptimality of delta hedging for Asian options. SIAM Journal on Financial Mathematics, 6(1): 352–385, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Kardaras:2010:MEM

[KP10a] Constantinos Kardaras and Eckhard Platen. Minimizing the expected market time to reach a certain wealth level. SIAM Journal on Financial Mathematics, 1 (1):16–29, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Kharroubi:2010:OPL

[KP10b] Idris Kharroubi and Huyên Pham. Optimal portfolio liquidation with execution cost and risk. SIAM Journal on Financial Mathematics, 1(1): 897–931, ???? 2010. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v1/i1/p897_s1.

Kramkov:2016:SAE

[KP16] Dmitry Kramkov and Sergio Pulido. Stability and analytic expansions of local solutions of systems of quadratic BS-DEs with applications to a price impact model. SIAM Journal on Financial Mathematics, 7(1): 567–587, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Kwak:2018:CPT

[KP18] Minsuk Kwak and Traian A.

Pirvu. Cumulative prospect theory with generalized hyperbolic skewed t distribution. SIAM Journal on Financial Mathematics, 9(1):54–89, ???? 2018. CO-DEN SJFMBJ. ISSN 1945-497X.

Kong:2022:RUP

[KQY22] Linghui Kong, Cong Qin, and Xingye Yue. Realization utility with path-dependent reference points. SIAM Journal on Financial Mathematics, 13(3): 1063–1111, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1411457.

Kratschmer:2010:ROS

[KS10] Volker Krätschmer and John Schoenmakers. Representations for optimal stopping under dynamic monetary utility functionals. SIAM Journal on Financial Mathematics, 1(1):811–832, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Kusnetsov:2019:ICF

[KV19] Michael Kusnetsov and Luitgard Anna Maria Veraart. Interbank clearing in financial networks with multiple maturities. SIAM Journal on Financial Mathematics, 10(1):37–67, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Kleisinger-Yu:2020:MPF

[KYKLR20] Xi Kleisinger-Yu, Vlatka Komaric, Martin Larsson, and Markus Regez. A multifactor polynomial framework for longterm electricity forwards with delivery period. SIAM Journal on Financial Mathematics, 11(3):928–957, ???? 2020. CO-DEN SJFMBJ. ISSN 1945-497X.

Lessy:2021:MML

[LDDD21] Djaffar Lessy, Nahla Dhib, Francine Diener, and Marc Diener. May microcredit lead to inclusion? SIAM Journal on Financial Mathematics, 12(3): 898–911, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Lelong:2018:DPA

[Lel18] Jérôme Lelong. Dual pricing of American options by Wiener chaos expansion. SIAM Journal on Financial Mathematics, 9(2): 493–519, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Lepinette:2016:RNA

[Lep16] Emmanuel Lepinette. Robust no arbitrage of the second kind with a continuum of assets and proportional transaction costs. SIAM Journal on Financial Mathematics, 7(1):104–123, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Levendorskii:2011:CPS

[Lev11] Sergei Levendorskii. Convergence of price and sensitivities in Carr’s randomization approximation globally and near barrier. SIAM Journal on Financial Mathematics, 2(1): 79–111, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p79_s1.

Levendorskii:2018:PAA

[Lev18] Sergei Levendorskii. Pricing arithmetic Asian options under Lévy models by backward induction in the dual space. SIAM Journal on Financial Mathematics, 9(1):1–27, ???? 2018. CO-DEN SJFMBJ. ISSN 1945-497X.

Liang:2010:OCR

[LHJ10] Jin Liang, Bei Hu, and Lishang Jiang. Optimal convergence rate of the binomial tree scheme for American options with jump diffusion and their free boundaries. SIAM Journal on Financial Mathematics, 1 (1):30–65, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Li:2022:HON

[Li22] Yunzhang Li. A high-order numerical method for BSPDEs with applications to mathematical finance. SIAM Journal on Financial Mathematics, 13(1): 147–178, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/20M1383252.

Liu:2015:PAV

[Liu15] Hsuan-Ku Liu. Properties of American volatility options in the mean-reverting 3/2 volatility model. SIAM Journal on Financial Mathematics, 6(1):53– 65, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Leung:2011:OTP

[LL11] Tim Leung and Mike Ludkovski. Optimal timing to purchase options. SIAM Journal on Financial Mathematics, 2 (1):768–793, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p768_s1.

Lorig:2016:VSD

[LLCMA16] Matthew Lorig, Oriol Lozano-Carbassé, and Rafael Mendoza-Arriaga. Variance swaps on defaultable assets and market implied time-changes. SIAM Journal on Financial Mathematics, 7 (1):273–307, ???? 2016. CO-DEN SJFMBJ. ISSN 1945-497X.

Li:2021:GTA

[LLL21] Juan Li, Wenqiang Li, and Gechun Liang. A game theoretical approach to homothetic robust forward investment performance processes in stochastic factor models. SIAM Journal on Financial Mathematics, 12(3): 867–897, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Landriault:2018:ESM

[LLLY18] David Landriault, Bin Li, Danping Li, and Virginia R. Young. Equilibrium strategies for the mean-variance investment problem over a random horizon. SIAM Journal on Financial Mathematics, 9(3):1046–1073, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Lillo:2023:ABL

[LLM^{+}23] Fabrizio Lillo, Giulia Livieri, Stefano Marmi, Anton Solomko, and Sandro Vaienti. Analysis of bank leverage via dynamical systems and deep neural networks. SIAM Journal on Financial Mathematics, 14(2): 598–643, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1412517.

Laruelle:2011:OSO

[LLP11] Sophie Laruelle, Charles-Albert Lehalle, and Gilles Pagès. Optimal split of orders across liquidity pools: a stochastic algorithm approach. SIAM Journal on Financial Mathematics, 2(1):1042–1076, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL http://epubs.
                                        siam.org/sifin/resource/1/
                                        sjfmbj/v2/i1/p1042_s1.

Landriault:2023:OSE

[LLP23] David Landriault, Bin Li, and José M. Pedraza. Optimal stopping for exponential Lévy models with weighted discounting. SIAM Journal on Financial Mathematics, 14(3):777– 811, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M1513538.

Li:2014:DBS

[LLS14] Xiao Li, Michael D. Lipkin, and Richard B. Sowers. Dynamics of bankrupt Stocks. SIAM Journal on Financial Mathematics, 5(1): 232–257, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Li:2019:ESA

[LLX19] Bin Li, Peng Luo, and Dewen Xiong. Equilibrium strategies for alpha-maxmin expected utility maximization. SIAM Journal on Financial Mathematics, 10(2):394–429, ???? 2019. CO-DEN SJFMBJ. ISSN 1945-497X.

Liu:2013:PSS

[LMK13] Ren Liu and Johannes Muhle-Karbe. Portfolio selection with small transaction costs and binding portfolio constraints. SIAM Journal on Financial Mathematics, 4(1):203–227, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Liebrich:2022:MUR

[LMS22] Felix-Benedikt Liebrich, Marco Maggis, and Gregor Svindland. Model uncertainty: a reverse approach. SIAM Journal on Financial Mathematics, 13(3): 1230–1269, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1425463.

Liebrich:2022:SVR

[LN22] Felix-Benedikt Liebrich and Max Nendel. Separability versus robustness of Orlicz spaces: Financial and economic perspectives. SIAM Journal on Financial Mathematics, 13(4):1344– 1378, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1418794.

Londono:2020:DEH

[Lon20] Jaime A. Londoño. Duesenberry equilibrium and heterogeneous agents. SIAM Journal on Financial Mathematics, 11(3): 659–689, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Lototsky:2021:KCS

[LP21] Sergey Lototsky and Austin Pollok. Kelly criterion: From a simple random walk to Lévy processes. SIAM Journal on Financial Mathematics, 12(1):342– 368, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Lopez:2021:EPJ

[LPY21] Dante Mata López, José Luis Pérez, and Kazutoshi Yamazaki. Effects of positive jumps of assets on endogenous bankruptcy and optimal capital structure: Continuous- and periodicobservation models. SIAM Journal on Financial Mathematics, 12(3):1112–1149, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Laachir:2016:BCM

[LR16] Ismail Laachir and Francesco Russo. BSDEs, Càdlàg martingale problems, and orthogonalization under basis risk. SIAM Journal on Financial Mathematics, 7(1):308–356, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Lorig:2016:POU

[LS16] Matthew Lorig and Ronnie Sircar. Portfolio optimization under local-stochastic volatility: Coefficient Taylor series approximations and implied Sharpe ratio. SIAM Journal on Financial Mathematics, 7(1):418–447, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Lin:2021:AOP

[LS21] Minglian Lin and Indranil Sen-Gupta. Analysis of optimal portfolio on finite and small-time horizons for a stochastic volatility market model. SIAM Journal on Financial Mathematics, 12(4):1596–1624, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Li:2018:WCR

[LSWY18] Lujun Li, Hui Shao, Ruodu Wang, and Jingping Yang. Worst-case range value-at-risk with partial information. SIAM Journal on Financial Mathematics, 9(1):190–218, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Lee:2020:RSF

[LSZ20] Junbeom Lee, Stephan Sturm, and Chao Zhou. A risksharing framework of bilateral contracts. SIAM Journal on Financial Mathematics, 11(2): 385–410, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Lamberton:2019:VFA

[LT19] Damien Lamberton and Giulia Terenzi. Variational formulation of American option prices in the Heston model. SIAM Journal on Financial Mathematics, 10(1):261–308, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Ludkovski:2011:SSG

[Lud11] Michael Ludkovski. Stochastic switching games and duopolistic competition in emissions markets. SIAM Journal on Financial Mathematics, 2(1): 488–511, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p488_s1.

Leung:2015:EVJ

[LW15] Tim Leung and Haohua Wan. ESO valuation with job termination risk and jumps in stock price. SIAM Journal on Financial Mathematics, 6(1):487–516, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Li:2015:AGM

[LX15] Cheng Li and Hao Xing. Asymptotic Glosten–Milgrom equilibrium. SIAM Journal on Financial Mathematics, 6(1):242–280, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Liang:2017:RHF

[LZ17] Gechun Liang and Thaleia Zariphopoulou. Representation of homothetic forward performance processes in stochastic factor models via ergodic and infinite horizon BSDE. SIAM Journal on Financial Mathematics, 8(1):

344–372, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Marchenko:2018:TCT

[MGH18] Ganna Marchenko, Patrick Gagliardini, and Illia Horenko. Towards a computationally tractable maximum entropy principle for nonstationary financial time series. SIAM Journal on Financial Mathematics, 9 (4):1249–1285, ???? 2018. CO-DEN SJFMBJ. ISSN 1945-497X.

Muhle-Karbe:2012:OPM

[MKPS12] Johannes Muhle-Karbe, Oliver Pfaffel, and Robert Stelzer. Option pricing in multivariate stochastic volatility models of OU type. SIAM Journal on Financial Mathematics, 3(1):66–94, ???? 2012. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v3/i1/p66_s1.

Martini:2022:NAS

[MM22] Claude Martini and Arianna Mingone. No arbitrage SVI. SIAM Journal on Financial Mathematics, 13(1):227–261, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X. URL https:/

/epubs.siam.org/doi/10.1137/
20M1351060.

Monnier:2013:RND

[Mon13a] Jean-Baptiste Monnier. Riskneutral density recovery via spectral analysis. SIAM Journal on Financial Mathematics, 4(1):

650–667, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Monoyios:2013:MCM

[Mon13b] Michael Monoyios. Malliavin calculus method for asymptotic expansion of dual control problems. SIAM Journal on Financial Mathematics, 4(1):884–915, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Mostovyi:2021:SIU

[Mos21]  Oleksii Mostovyi.  Stability

of the indirect utility process. SIAM Journal on Financial Mathematics, 12(2):641– 671, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Marisu:2023:BEO

[MP23] Godeliva Petrina Marisu and Chi Seng Pun. Bayesian estimation and optimization for learning sequential regularized portfolios. SIAM Journal on Financial Mathematics, 14(1): 127–157, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1427176.

Montes:2014:MCV

[MPR14] Juan Miguel Montes, Valentina Prezioso, and Wolfgang J. Runggaldier. Monte Carlo variance reduction by conditioning for pricing with underlying a continuous-time finite state Markov process. SIAM Journal on Financial Mathematics, 5(1): 557–580, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Mercuri:2021:FMA

[MPR21] Lorenzo Mercuri, Andrea Perchiazzo, and Edit Rroji. Finite mixture approximation of CARMA(p, q) models. SIAM Journal on Financial Mathematics, 12(4):1416–1458, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Mastrolia:2018:PAP

[MR18] Thibaut Mastrolia and Zhenjie Ren. Principal-agent problem with common agency without communication. SIAM Journal on Financial Mathematics, 9(2): 775–799, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Molchanov:2010:MEP

[MS10] Ilya Molchanov and Michael Schmutz. Multivariate extension of put-call symmetry. SIAM Journal on Financial Mathematics, 1(1):396–426, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Milstein:2014:CMS

[MS14] G. N. Milstein and V. Spokoiny.

Construction of mean-self-financing strategies for European options under regime-switching. SIAM Journal on Financial Mathematics, 5(1):532–556, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Malham:2021:SED

[MSW21] Simon J. A. Malham, Jiaqi Shen, and Anke Wiese. Series expansions and direct inversion for the Heston model. SIAM Journal on Financial Mathematics, 12(1): 487–549, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Mania:2017:RPD

[MT17] Michael Mania and Revaz Tevzadze. On regularity of primal and dual dynamic value functions related to investment problems and their representations as backward stochastic PDE solutions. SIAM Journal on Financial Mathematics, 8(1): 483–503, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Mocha:2013:SCU

[MW13] Markus Mocha and Nicholas Westray. The stability of the constrained utility maximization problem: a BSDE approach. SIAM Journal on Financial Mathematics, 4(1):117– 150, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Mao:2020:RAR

[MW20] Tiantian Mao and Ruodu Wang. Risk aversion in regulatory capital principles. SIAM Journal on Financial Mathematics, 11(1): 169–200, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Meng:2022:ODR

[MWZZ22] Hui Meng, Pengyu Wei, Wanlu Zhang, and Sheng Chao Zhuang. Optimal dynamic reinsurance under heterogeneous beliefs and CARA utility. SIAM Journal on Financial Mathematics, 13(3): 903–943, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1411093.

Musiela:2010:PCU

[MZ10] M. Musiela and T. Zariphopoulou.

Portfolio choice under spacetime monotone performance criteria. SIAM Journal on Financial Mathematics, 1(1):326–365, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Ning:2023:AFI

[NJZB23] Brian (Xin) Ning, Sebastian Jaimungal, Xiaorong Zhang, and Maxime Bergeron. Arbitragefree implied volatility surface generation with variational autoencoders. SIAM Journal on Financial Mathematics, 14(4): 1004–1027, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1443546.

Nicole:2013:ECB

[NM13] El Karoui Nicole and Mrad Mohamed. An exact connection between two solvable SDEs and a nonlinear utility stochastic PDE. SIAM Journal on Financial Mathematics, 4(1):697–736, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Nicolato:2017:IJD

[NPS17] E. Nicolato, C. Pisani, and D. Sloth. The impact of jump distributions on the implied volatility of variance. SIAM Journal on Financial Mathematics, 8(1):28–53, ???? 2017. CO-DEN SJFMBJ. ISSN 1945-497X.

Neufeld:2021:MFP

[NS21] Ariel Neufeld and Julian Sester. Model-free price bounds under dynamic option trading. SIAM Journal on Financial Mathematics, 12(4):1307–1339, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Neuman:2022:OSA

[NV22] Eyal Neuman and Moritz Voß. Optimal signal-adaptive trading with temporary and transient price impact. SIAM Journal on Financial Mathematics, 13(2): 551–575, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/20M1375486.

Ning:2021:WPS

[NW21] Ning Ning and Jing Wu. Wellposedness and stability analysis of two classes of generalized stochastic volatility models. SIAM Journal on Financial Mathematics, 12(1):79–109, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Nadtochiy:2013:ASS

[NZ13] Sergey Nadtochiy and Thaleia Zariphopoulou. An approximation Scheme for solution to the optimal investment problem in incomplete markets. SIAM Journal on Financial Mathematics, 4(1):494–538, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Nadtochiy:2019:OCF

[NZ19] Sergey Nadtochiy and Thaleia Zariphopoulou. Optimal contract for a fund manager with capital injections and endogenous trading constraints. SIAM Journal on Financial Mathematics, 10(3):698–722, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Nutz:2022:RDR

[NZ22] Marcel Nutz and Yuchong Zhang. Reward design in risktaking contests. SIAM Journal on Financial Mathematics, 13 (1):129–146, ???? 2022. CO-DEN SJFMBJ. ISSN 1945-

497X. URL https://epubs.
siam.org/doi/10.1137/21M1397386.

Ogetbil:2023:EDF

[ÖH23] Orcan Ögetbil and Bernhard Hientzsch. Extensions of dupire formula: Stochastic interest rates and stochastic local volatility. SIAM Journal on Financial Mathematics, 14(2):452– 474, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1390906.

Papanicolaou:2013:DRD

[Pap13] Andrew Papanicolaou. Dimension reduction in discrete time portfolio optimization with partial information. SIAM Journal on Financial Mathematics, 4(1): 916–960, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Papanicolaou:2018:ESC

[Pap18]  A. Papanicolaou.  Extreme-

strike comparisons and structural bounds for SPX and VIX options. SIAM Journal on Financial Mathematics, 9(2):401– 434, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Putzig:2010:OAF

[PBH10] L. Putzig, D. Becherer, and I. Horenko. Optimal allocation of a futures portfolio utilizing numerical market phase detection. SIAM Journal on Financial Mathematics, 1(1):752–779, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Park:2010:PSU

[PO10] Sungwoo Park and Dianne P. O’Leary. Portfolio selection using Tikhonov filtering to estimate the covariance matrix. SIAM Journal on Financial Mathematics, 1(1):932– 961, ???? 2010. CO-DEN SJFMBJ. ISSN 1945-

                                        497X. URL http://epubs.
                                        siam.org/sifin/resource/1/
                                        sjfmbj/v1/i1/p932_s1.

Pennanen:2010:HCP

[PP10] Teemu Pennanen and Irina Penner. Hedging of claims with physical delivery under convex transaction costs. SIAM Journal on Financial Mathematics, 1(1): 158–178, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Pagliarani:2013:AEL

[PPR13] Stefano Pagliarani, Andrea Pascucci, and Candia Riga. Adjoint expansions in local Lévy models. SIAM Journal on Financial Mathematics, 4(1):265–296, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Pages:2018:PAA

[PPS18] Gilles Pagès, Olivier Pironneau, and Guillaume Sall. The parareal algorithm for American options. SIAM Journal on Financial Mathematics, 9(3): 966–993, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Predoiu:2011:OEG

[PSS11] Silviu Predoiu, Gennady Shaikhet, and Steven Shreve. Optimal execution in a general one-sided limit-order book. SIAM Journal on Financial Mathematics, 2(1):183–212, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v2/i1/p183_s1.

Pun:2021:SLA

[Pun21]  Chi Seng Pun.  A sparse

learning approach to relativevolatility-managed portfolio selection. SIAM Journal on Financial Mathematics, 12(1):410– 445, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Pierre:2017:NAC

[PVW17] Erwan Pierre, Stéphane Villeneuve, and Xavier Warin. Numerical approximation of a cashconstrained firm value with investment opportunities. SIAM Journal on Financial Mathematics, 8(1):54–81, ???? 2017. CO-DEN SJFMBJ. ISSN 1945-497X.

Pun:2016:RDM

[PW16] Chi Seng Pun and Hoi Ying Wong. Resolution of degeneracy in Merton’s portfolio problem. SIAM Journal on Financial Mathematics, 7(1):786–811, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Park:2022:RCI

[PW22] Kyunghyun Park and Hoi Ying Wong. Robust consumptioninvestment with return ambiguity: a dual approach with volatility ambiguity. SIAM Journal on Financial Mathematics, 13(3):802–843, ???? 2022. CODEN SJFMBJ. ISSN 1945-

497X. URL https://epubs.
siam.org/doi/10.1137/21M1440189.

Pirjol:2016:SMA

[PZ16] Dan Pirjol and Lingjiong Zhu. Short maturity Asian options in local volatility models. SIAM Journal on Financial Mathematics, 7(1):947–992, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Qin:2019:BGP

[QC19] Cong Qin and Xinfu Chen. On balanced growth path solutions of a knowledge diffusion and growth model. SIAM Journal on Financial Mathematics, 10(1):

130–155, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Qi:2023:GCM

[Qi23]  Hou-Duo Qi.  Geometric

characterization of maximum diversification return portfolio via Rao’s quadratic entropy. SIAM Journal on Financial Mathematics, 14(2):525– 556, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/22M1492313.

Rasonyi:2015:OIN

[Rás15] Miklós Rásonyi. Optimal investment with nonconcave utilities in discrete-time markets. SIAM Journal on Financial Mathematics, 6(1):517–529, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Redmann:2021:LDA

[RBG21] Martin Redmann, Christian Bayer, and Pawan Goyal. Lowdimensional approximations of high-dimensional asset price models. SIAM Journal on Financial Mathematics, 12(1): 1–28, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Risk:2018:SDS

[RL18] Jimmy Risk and Michael Ludkovski. Sequential design and spatial modeling for portfolio tail risk measurement. SIAM Journal on Financial Mathematics, 9(4):1137–1174, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Rogers:2010:DVH

[Rog10] L. C. G. Rogers. Dual valuation and hedging of Bermudan options. SIAM Journal on Financial Mathematics, 1(1):604–608, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Robert:2010:MHE

[RR10] Christian Y. Robert and Mathieu Rosenbaum. On the microstructural hedging error. SIAM Journal on Financial Mathematics, 1(1):427–453, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.

Rodriguez:2012:EAD

[RS12] J. Orozco Rodriguez and F. Santosa. Estimation of asset distributions from option prices: Analysis and regularization. SIAM Journal on Financial Mathematics, 3(1):374–401, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Rheinlander:2014:QSD

[RS14] Thorsten Rheinländer and Michael Schmutz. Quasi–self-dual exponential Lévy processes. SIAM Journal on Financial Mathematics, 5(1):656–684, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Richter:2017:DTT

[RT17] Anja Richter and Josef Teichmann. Discrete time term structure theory and consistent recalibration models. SIAM Journal on Financial Mathematics, 8(1): 504–531, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Richard:2023:DTS

[RTY23] Alexandre Richard, Xiaolu Tan, and Fan Yang. On the discretetime simulation of the rough Heston model. SIAM Journal on Financial Mathematics, 14(1): 223–249, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1443807.

Reisinger:2012:UPI

[RW12] C. Reisinger and J. H. Witte. On the use of policy iteration as an easy way of pricing American options. SIAM Journal on Financial Mathematics, 3(1):459–478, ???? 2012. CODEN SJFMBJ. ISSN 1945-497X.

Robertson:2017:LTO

[RX17] Scott Robertson and Hao Xing.

Long-term optimal investment in matrix valued factor models. SIAM Journal on Financial Mathematics, 8(1):400–434, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Ruf:2020:IPT

[RX20] Johannes Ruf and Kangjianan Xie. The impact of proportional transaction costs on systematically generated portfolios. SIAM Journal on Financial Mathematics, 11(3):881–896, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Sabino:2023:NTS

[Sab23] Piergiacomo Sabino. Normal tempered stable processes and the pricing of energy derivatives. SIAM Journal on Financial Mathematics, 14(1):99– 126, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1425207.

Saporito:2020:SCP

[Sap20]  Yuri F. Saporito.  Short

communication: Pricing pathdependent derivatives under multiscale stochastic volatility models: a Malliavin representation. SIAM Journal on Financial Mathematics, 11(3):SC– 14–SC–25, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Schoneborn:2015:OTE

[Sch15]  Torsten Schöneborn.  Opti-

mal trade execution for timeinconsistent mean-variance criteria and risk functions. SIAM Journal on Financial Mathematics, 6(1):1044–1067, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Sekine:2013:LTO

[Sek13] Jun Sekine. Long-term optimal investment with a generalized drawdown constraint. SIAM Journal on Financial Mathematics, 4(1):452–473, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Shinozaki:2017:CTO

[Shi17] Yuji Shinozaki. Construction of a third-order K-scheme and its application to financial models. SIAM Journal on Financial Mathematics, 8(1):901–932, ????

2017. CODEN SJFMBJ. ISSN 1945-497X.

Spiliopoulos:2015:DCL

[SS15] Konstantinos Spiliopoulos and Richard B. Sowers. Default clustering in large pools: Large deviations. SIAM Journal on Financial Mathematics, 6(1):86–116, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Sirignano:2017:SGD

[SS17] Justin Sirignano and Konstantinos Spiliopoulos. Stochastic gradient descent in continuous time. SIAM Journal on Financial Mathematics, 8(1):933–961, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Schatz:2019:NIM

[SS19] Michael Schatz and Didier Sornette. A nonuniformly integrable martingale bubble with a crash. SIAM Journal on Financial Mathematics, 10(2):615– 631, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Schied:2018:MFP

[SSV18] Alexander Schied, Leo Speiser, and Iryna Voloshchenko. Modelfree portfolio theory and its functional master formula. SIAM Journal on Financial Mathematics, 9(3):1074–1101, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Shkolnikov:2016:AAF

[SSZ16] Mykhaylo Shkolnikov, Ronnie Sircar, and Thaleia Zariphopoulou. Asymptotic analysis of forward performance processes in incomplete markets and their ill-posed HJB equations. SIAM Journal on Financial Mathematics, 7(1):588–618, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Schied:2017:HFL

[SSZ17] Alexander Schied, Elias Strehle, and Tao Zhang. High-frequency limit of Nash equilibria in a market impact game with transient price impact. SIAM Journal on Financial Mathematics, 8(1): 589–634, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Strong:2014:GFG

[Str14] Winslow Strong. Generalizations of functionally generated portfolios with applications to statistical arbitrage. SIAM Journal on Financial Mathematics, 5(1):472–492, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Swishchuk:2017:SMM

[SV17] Anatoliy Swishchuk and Nelson Vadori. A semi-Markovian modeling of limit order markets. SIAM Journal on Financial Mathematics, 8(1):240–273, ???? 2017. CODEN SJFMBJ. ISSN 1945-497X.

Shreve:2022:EC

[SW22] Steven Shreve and Jing Wang. Escrow and clawback. SIAM Journal on Financial Mathematics, 13(3):1191–1229, ???? 2022. CODEN SJFMBJ. ISSN 1945-

497X. URL https://epubs.
siam.org/doi/10.1137/21M1455619.

Saporito:2021:PDD

[SZ21] Yuri F. Saporito and Zhaoyu Zhang. Path-dependent deep Galerkin method: a neural network approach to solve pathdependent partial differential equations. SIAM Journal on Financial Mathematics, 12(3): 912–940, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Shen:2022:MVP

[SZ22a] Yang Shen and Bin Zou.

Mean-variance portfolio selection in contagious markets. SIAM Journal on Financial Mathematics, 13(2):391– 425, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/20M1320560.

Shen:2022:SCC

[SZ22b] Yang Shen and Bin Zou.

Short communication: Coneconstrained monotone meanvariance portfolio selection under diffusion models. SIAM Journal on Financial Mathematics, 13(4):SC99–SC112, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X. URL https://

epubs.siam.org/doi/10.1137/
22M1487527.

Schoenmakers:2013:ODM

[SZH13] John Schoenmakers, Jianing Zhang, and Junbo Huang. Optimal dual martingales, their analysis, and application to new algorithms for Bermudan products. SIAM Journal on Financial Mathematics, 4(1):86–116, ???? 2013. CODEN SJFMBJ. ISSN 1945-497X.

Trevinno-Aguilar:2015:DPS

[TA15] Erick Trevinño-Aguilar. Duality in a problem of static partial hedging under convex constraints. SIAM Journal on Financial Mathematics, 6(1):1152– 1170, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Tissot-Daguette:2022:SCP

[TD22] Valentin Tissot-Daguette. Short communication: Projection of functionals and fast pricing of exotic options. SIAM Journal on Financial Mathematics, 13(2):SC74–SC86, ???? 2022. CODEN SJFMBJ. ISSN 1945-

497X. URL https://epubs.
siam.org/doi/10.1137/21M1451439.

Tehranchi:2016:UBB

[Teh16] Michael R. Tehranchi. Uniform bounds for Black–Scholes implied volatility. SIAM Journal on Financial Mathematics, 7(1): 893–916, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Tian:2023:PPT

[Tia23] Dejian Tian. Pricing principle via Tsallis relative entropy in incomplete markets. SIAM Journal on Financial Mathematics, 14(1):250– 278, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/22M1471614.

Tan:2018:OTP

[TT18] Zongjun Tan and Peter Tankov.

Optimal trading policies for wind energy producer. SIAM Journal on Financial Mathematics, 9(1):315–346, ???? 2018. CODEN SJFMBJ. ISSN 1945-497X.

Tsang:2020:DLS

[TW20] Ka Ho Tsang and Hoi Ying Wong. Deep-learning solution to portfolio selection with serially dependent returns. SIAM Journal on Financial Mathematics, 11(2):593–619, ???? 2020. CODEN SJFMBJ. ISSN 1945-497X.

Takahashi:2012:AEP

[TY12] Akihiko Takahashi and Toshihiro Yamada. An asymptotic expansion with push-down of Malliavin weights. SIAM Journal on Financial Mathematics, 3(1):95–136, ???? 2012. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.
siam.org/sifin/resource/1/
sjfmbj/v3/i1/p95_s1.

Veraguas:2016:RUM

[VF16] Julio D. Backhoff Veraguas and Joaquín Fontbona. Robust utility maximization without model compactness. SIAM Journal on Financial Mathematics, 7(1): 70–103, ???? 2016. CODEN SJFMBJ. ISSN 1945-497X.

Staden:2023:BBD

[VFL23] Pieter M. Van Staden, Peter A. Forsyth, and Yuying Li. Beating a benchmark: Dynamic programming may not be the right numerical approach. SIAM Journal on Financial Mathematics, 14(2):407– 451, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/22M1530070.

Vellekoop:2022:ESA

[VG22] Michel Vellekoop and Marcellino Gaudenzi. Exact solutions and approximations for optimal investment strategies and indifference prices. SIAM Journal on Financial Mathematics, 13(2): 491–520, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1393303.

Vigna:2022:TOP

[Vig22] Elena Vigna. Tail optimality and preferences consistency for intertemporal optimization problems. SIAM Journal on Financial Mathematics, 13(1): 295–320, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1435422.

Veraguas:2022:SCO

[VRT22] Julio Backhoff Veraguas, A. Max Reppen, and Ludovic Tangpi. Stochastic control of optimized certainty equivalents. SIAM Journal on Financial Mathematics, 13(3):745–772, ???? 2022. CODEN SJFMBJ. ISSN 1945-

497X. URL https://epubs.
siam.org/doi/10.1137/21M1407732.

vanStaden:2019:MQV

[vSDF19] Pieter M. van Staden, Duy-Minh Dang, and Peter A. Forsyth. Mean-quadratic variation portfolio optimization: a desirable alternative to timeconsistent mean-variance optimization? SIAM Journal on Financial Mathematics, 10(3): 815–856, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

vanStaden:2021:DTW

[vSDF21] Pieter M. van Staden, Duy-Minh Dang, and Peter A. Forsyth. On the distribution of terminal wealth under dynamic mean-variance optimal investment strategies. SIAM Journal on Financial Mathematics, 12(2):566–603, ???? 2021. CO-DEN SJFMBJ. ISSN 1945-497X.

Wang:2022:PFS

[Wan22]  Gu Wang.  Performance

fees with stochastic benchmark. SIAM Journal on Financial Mathematics, 13(2):619– 652, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

URL https://epubs.siam.
org/doi/10.1137/21M1401826.

Ware:2013:ASL

[War13] Antony Ware. Accurate semi-Lagrangian time stepping for stochastic optimal control problems with application to the valuation of natural gas storage. SIAM Journal on Financial Mathematics, 4(1):427–451, ????

2013. CODEN SJFMBJ. ISSN 1945-497X.

Wang:2015:HSC

[WBT15] Ruodu Wang, Valeria Bignozzi, and Andreas Tsanakas. How superadditive can a risk measure be? SIAM Journal on Financial Mathematics, 6(1):776–803, ???? 2015. CODEN SJFMBJ. ISSN 1945-497X.

Wang:2021:EUM

[WX21] Xiangyu Wang and Jianming Xia. Expected utility maximization with stochastic dominance constraints in complete markets. SIAM Journal on Financial Mathematics, 12(3): 1054–1111, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Wang:2022:SCM

[WXXY22] Xiangyu Wang, Jianming Xia, Zuo Quan Xu, and Zhou Yang. Short communication: Minimal quantile functions subject to stochastic dominance constraints. SIAM Journal on Financial Mathematics, 13(3): SC87–SC98, ???? 2022. CO-DEN SJFMBJ. ISSN 1945-

497X. URL https://epubs.
siam.org/doi/10.1137/22M1488557.

Wu:2011:MRS

[WYZ11] Zhijian Wu, Chunhui Yu, and Xiaohua Zheng. Managing risk with short-term futures contracts. SIAM Journal on Financial Mathematics, 2(1): 715–726, ???? 2011. CO-DEN SJFMBJ. ISSN 1945-

497X. URL http://epubs.

siam.org/sifin/resource/1/
sjfmbj/v2/i1/p715_s1.

Yamada:2022:SCG

[Yam22]  Toshihiro Yamada.  Short

communication: A Gaussian Kusuoka approximation without solving random ODEs. SIAM Journal on Financial Mathematics, 13(1):SC1–SC11, ???? 2022. CODEN SJFMBJ. ISSN 1945-

497X. URL https://epubs.
siam.org/doi/10.1137/21M1433915.

Yang:2023:RCP

[YZZ23] Zhou Yang, Jing Zhang, and Chao Zhou. Robust control problems of BSDEs coupled with value functions. SIAM Journal on Financial Mathematics, 14 (3):721–750, ???? 2023. CO-DEN SJFMBJ. ISSN 1945-

497X. URL https://epubs.
siam.org/doi/10.1137/22M1511977.

Zhang:2016:CNP

[ZCLG16] Geliang Zhang, Hugh Christensen, Guolong Li, and Simon Godsill. A correction note for price dynamics in a Markovian limit order market. SIAM Journal on Financial Mathematics, 7 (1):152–158, ???? 2016. CO-DEN SJFMBJ. ISSN 1945-497X.

Zhang:2023:SCS

[Zha23] Jianfeng Zhang. Short communication: Is a sophisticated agent always a wise one? SIAM Journal on Financial Mathematics, 14(4):SC42– SC48, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/23M1569137.

Zhitlukhin:2023:CGS

[Zhi23] Mikhail Zhitlukhin. Capital growth and survival strategies in a market with endogenous prices. SIAM Journal on Financial Mathematics, 14(3):812– 837, ???? 2023. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1394370.

Zhou:2021:UMW

[Zho21] Zhou Zhou. Utility maximization when shorting American options. SIAM Journal on Financial Mathematics, 12(1):47–78, ???? 2021. CODEN SJFMBJ. ISSN 1945-497X.

Zhang:2022:AMC

[ZL22] Gongqiu Zhang and Lingfei Li. Analysis of Markov chain approximation for diffusion models with nonsmooth coefficients. SIAM Journal on Financial Mathematics, 13(3):1144– 1190, ???? 2022. CODEN SJFMBJ. ISSN 1945-497X.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1440098.

Zhang:2013:EPE

[ZO13] B. Zhang and C. W. Oosterlee. Efficient pricing of Europeanstyle Asian options under exponential Lévy processes based on Fourier cosine expansions. SIAM Journal on Financial Mathematics, 4(1):399–426, ???? 2013.

CODEN SJFMBJ. ISSN 1945-497X.

Zheng:2014:MBA

[ZRA14] Ban Zheng, François Roueff, and Frédéric Abergel. Modelling bid and ask prices using constrained Hawkes processes: Ergodicity and scaling limit. SIAM Journal on Financial Mathematics, 5 (1):99–136, ???? 2014. CODEN SJFMBJ. ISSN 1945-497X.

Zeng:2019:MMC

[ZX19] Ailing Zeng and Jungong Xue.

Multilevel Monte Carlo method for path-dependent barrier interest rate derivatives. SIAM Journal on Financial Mathematics, 10(1):214–242, ???? 2019. CODEN SJFMBJ. ISSN 1945-497X.

Zariphopoulou:2010:MIR

[ZŽ10] Thaleia Zariphopoulou and Gordan Žitković. Maturityindependent risk measures. SIAM Journal on Financial Mathematics, 1(1):266–288, ???? 2010. CODEN SJFMBJ. ISSN 1945-497X.